### v2

# GraphOracle: Efficient Fully-Inductive Knowledge Graph Reasoning via Relation-Dependency Graphs

Abstract
Knowledge graph reasoning in the fully-inductive setting—where both entities and relations at test time are unseen during training—remains an open challenge. In this work, we introduce GraphOracle, a novel framework that achieves robust fully-inductive reasoning by transforming each knowledge graph into a Relation-Dependency Graph (RDG). The RDG encodes directed precedence links between relations, capturing essential compositional patterns while drastically reducing graph density. Conditioned on a query relation, a multi-head attention mechanism propagates information over the RDG to produce context-aware relation embeddings. These embeddings then guide a second GNN to perform inductive message passing over the original knowledge graph, enabling prediction on entirely new entities and relations. Comprehensive experiments on 60 benchmarks demonstrate that GraphOracle outperforms prior methods by up to 25% in fully-inductive and 28% in cross-domain scenarios. Our analysis further confirms that the compact RDG structure and attention-based propagation are key to efficient and accurate generalization
1 Introduction
Knowledge graphs (KGs) encode structured knowledge as entity–relation–entity triples, serving as the backbone for scientific discovery, web-scale reasoning, and intelligent systems (Cai2025FusionKGLLM; Bai2025AutoSchemaKG; Dong2025GLMTripleGen; guo2025decouplingcontinualsemanticsegmentation; lin2025seagentselfevolutiontrajectoryoptimization; du2025graphmaster). The central challenge in KG reasoning is link prediction: given an incomplete graph and a query $(h,r,?)$, predict the missing tail entity $t$ (Du2025Mixture). Inductive KG reasoning requires models to generalize to facts that were not explicitly observed during training. In the most challenging scenarios, models must perform fully-inductive (zhang2025trix; cui2024promptkg) reasoning—handling both unseen entities and relations during inference. Cross-domain generalization (wang2025mdgfm; hong2025stability; lin2025unifiedgnn) further requires reasoning over entirely different knowledge graphs containing 100% novel entities and relations. Both settings pose a fundamental compositional generalization challenge: models must recombine learned relational patterns to infer new facts in completely unfamiliar contexts.
Recent approaches to fully-inductive reasoning, including INGRAM (lee2023ingram) and ULTRA (galkin2024ultra), attempt to address this challenge by constructing auxiliary relation graphs $G_{R}=(R,E_{R})$ that capture transferable relational patterns. However, as knowledge graphs scale, this paradigm exposes two fundamental limitations that severely hinder their effectiveness in real-world applications. (1) These methods rely on co-occurrence statistics to connect relations, creating dense graphs with $|E_{R}|=\Theta(|R|^{2})$ edges that inflate computational costs to $O(|R|^{3}\cdot H)$ for $H$-layer message passing. The proliferation of spurious connections obscures meaningful compositional signals, while symmetric treatment of relation pairs erases the inherent directionality of logical composition. (2) Current models compress each relation into a single fixed embedding, forcing individual representations to capture diverse semantic roles across vastly different query contexts. For instance, the relation ”associated with” may connect proteins to diseases in biomedical contexts but link authors to topics in academic graphs, yet existing methods use the same representation regardless of context.
These observations raise a critical research question: How to design a KG reasoning framework that captures meaningful relational dependencies while maintaining computational efficiency? This requires moving beyond dense co-occurrence graphs to a sparse, directed structure that preserves compositional patterns, and replacing fixed relation embeddings with dynamic, query-dependent representations.
To realize this design, we propose GraphOracle, a relation-centric framework that transforms entity–relation interactions into a compact Relation-Dependency Graph (RDG). Unlike INGRAM and ULTRA, which generate excessive connections, our RDG retains only meaningful directed precedence links, yielding significantly fewer edges yet richer compositional patterns, with its effectiveness remaining stable as the number of relations increases. To capture relation dependencies for specific query, we design a multi-head attention mechanism that recursively propagates information over the RDG, dynamically assembling relation recipes conditioned on the query. By pre-training on four KGs in the general domain, GraphOracle only needs minimal finetune to achieve exceptional adaptability across transductive, inductive, and cross-domain reasoning tasks, improving performance by over 16.8% on average compared to state-of-the-art methods. Our key contributions in this work can be summarized as follows:
  • We introduce GraphOracle, a relation-centric foundation model that converts knowledge graphs into RDGs, explicitly encoding compositional patterns while reducing the number of edges on the relation graph compared to prior approaches.
  • We develop a query-dependent multi-head attention mechanism that dynamically propagates information over the RDG, yielding domain-invariant relation embeddings that enable generalization to unseen entities, relations, and graphs.
  • Extensive experiments across 60 benchmarks show that GraphOracle consistently outperforms SOTA methods, with particularly strong results in both fully-inductive and cross-domain settings, demonstrating its robustness and generalization capability in challenging scenarios.
2 Related Works
Knowledge Graph Reasoning
A Knowledge Graph (KG) consists of sets of entities $\mathcal{V}$, relations $\mathcal{R}$, and fact triples $\mathcal{F}\subseteq(\mathcal{V}\times\mathcal{R}\times\mathcal{V})$ as $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{F})$. $(e_{q},r_{q},e_{a})$ is a triple in KG where $e_{q},e_{a}\in\mathcal{V}$ and $r_{q}\in\mathcal{R}$. Knowledge graph reasoning encompasses several increasingly challenging settings based on what information is available during training versus inference. In the transductive setting, both entities and relations remain fixed: $(\mathcal{V}_{tra}=\mathcal{V}_{inf})\land(\mathcal{R}_{tra}=\mathcal{R}_{inf})$. This allows models to learn fixed embeddings for all components. The entity-inductive setting introduces unseen entities at inference while keeping relations fixed: $(\mathcal{V}_{tra}\neq\mathcal{V}_{inf})\land(\mathcal{R}_{tra}=\mathcal{R}_{inf})$. Most challenging is the fully-inductive setting where both entities and relations are novel: $(\mathcal{V}_{tra}\neq\mathcal{V}_{inf})\land(\mathcal{R}_{tra}\neq\mathcal{R}_{inf})$. Beyond these, cross-domain reasoning requires transferring to entirely different knowledge graphs with no shared entities or relations, demanding the most robust generalization capabilities.
2.1 Transductive Reasoning
Transductive methods assume all entities and relations at inference have been seen during training, enabling the use of fixed relation embeddings. Models like ConvE (ConvE), RotatE (RotatE) and DuASE (li2024duase) learn low-dimensional relation and entity embeddings directly, while GNN variants such as R-GCN (schlichtkrull2018modeling) implement relation-specific message passing that effectively parameterizing relation influence via learned embedding-like transformations. These embedding-based approaches form strong baselines but fundamentally cannot generalize beyond their training vocabulary.
2.2 Entity Inductive Reasoning
Entity inductive KG reasoning relaxes the entity constraint while maintaining fixed relation embeddings. Early solutions leveraged auxiliary cues—text descriptions in content-masking models (shi2018open) or ontological features in OntoZSL (geng2021ontozsl)—and symbolic rule learners such as AMIE (galarraga2013amie) and NeuralLP (yang2017neural). More recent approaches like DRUM (DRUM) employ differentiable rule chaining, while RLogic (RLogic) uses symbolic rule matching. SOTA GNN-based methods including GraIL (teru2020inductive), PathCon (wang2021pathcon), NBFNet (zhu2021nbfnet), RED-GNN (Zhang2022RelationalDigraph), A*Net (Zhu2023AStarNet) and AdaProp (Zhang2023AdaProp) propagate messages along relational paths to accommodate new entities—yet they still rely on fixed relation embeddings, limiting their applicability to scenarios with novel relations.
2.3 Fully-Inductive Reasoning
Fully-inductive settings demand handling both unseen entities and relations, requiring explicit relation graph structures. RMPI (geng2023relational) and INGRAM (lee2023ingram) pioneer this direction by constructing undirected relation graphs; however, RMPI is limited to subgraph extraction, and INGRAM’s degree discretization hampers transfer across graphs with different relation distributions. ISDEA (gao2023double) and MTDEA (zhou2023ood) adopt double-equivariant GNNs, but their computational overhead restricts scalability. ULTRA (galkin2024ultra) advances this with interaction-conditioned relation graphs that adapt based on query context. TRIX (zhang2025trix) introduces expressive adjacency motifs for richer relation modeling, while KG-ICL (cui2024kgicl) employs prompt-based relation graphs.
Method Ent-Ind. Full-Ind. Cross-Dom. Relation Representation
RotatE & DuASE Relation Embedding
A*Net & AdaProp Relation Embedding
DRUM Differentiable rule chaining
RLogic Symbolic rule matching
INGRAM Undirected RG
ULTRA Interaction-Conditioned RG
TRIX Expressive Adjacency Motifs RG
KG-ICL Prompt RG
GraphOracle Relation-Dependency Graph
Table 1: Comparison of inductive capabilities and relation representations. “RG” is short for “relation graph”.
2.4 Cross-domain Reasoning
Cross-domain KG reasoning represents the frontier of generalization, transferring patterns to graphs with entirely new entities and relations. Early work relied on domain-agnostic logical rules; recent advances leverage pre-trained graph foundation models. MDGFM (wang2025mdgfm) introduces multi-domain contrastive pre-training, while SAMGPT (zhang2025samgpt) demonstrates strong transfer without textual signals. Stability-GNN (hong2025stability) addresses structural shift through adversarial perturbations, and UnifiedGNN (lin2025unifiedgnn) jointly handles multiple inductive settings via relation adapters. RiemannGFM (liu2025riemanngfm) incorporates geometric regularization, while Text-Free MDGPT (li2025mfgpt) and GraphMFM (cheng2024graphmfm) scale cross-domain pre-training through modality-agnostic masked modeling.
As summarized in Table 1, KG reasoning methods progress from fixed relation embeddings (transductive) to rule-based reasoning (entity-inductive) to explicit relation graphs (fully-inductive). While existing fully-inductive methods construct undirected or interaction-conditioned graphs, they remain limited to single-domain scenarios. GraphOracle uniquely introduces directed relation-dependency graphs that capture compositional patterns, enabling the first successful cross-domain generalization.
Figure 1: Overview of the GraphOracle process that predicts the answer entity $e_{a}$ from a given query $(e_{1},r_{1},?)$: Given a Knowledge Graph, we first construct the Relation Dependency Graph (RDG). Then, a multi-head attention mechanism combined with a GNN is used to propagate messages among RDG to obtain relation representations $\bm{h}_{r|r_{q}}^{L_{r}}$, which are then used in another GNN for message passing over entity representations. Finally, the candidate entities are scored and evaluated based on the aggregated entity representations, and then ranked for answer entity prediction.
3 Preliminary
Given a query with missing answer $(e_{q},r_{q},?)$, the goal is to find an answer entity $e_{a}$ such that $(e_{q},r_{q},e_{a})$ is true. Most state-of-the-art models leverage GNN to aggregate relational paths and can be formulated as the following recursive function, where each candidate entity $e_{y}$ at step $\ell$ accumulates information from its in-neighbors:
$\bm{h}_{r_{q}}^{\ell}(e_{q},e_{y})=\bigoplus_{(e_{x},r,e_{y})\in\mathcal{N}(e_{y})}\!\!\bm{h}_{r_{q}}^{\ell-1}(e_{q},e_{x})\otimes\phi(r,r_{q}),$ (1)
where $\bm{h}_{r_{q}}^{0}(e_{q},e)=\bm{1}$ if $e=e_{q}$, and $\bm{0}$ otherwise, $\oplus/\otimes$ are learnable additive and multiplicative operators. $\phi$ encodes relation-type compatibility. After $L$ steps’ iteration, the answer is ranked by the score $s(e_{q},r_{q},e_{a})=\bm{w}_{s}^{\top}\bm{h}_{r_{q}}^{L}(e_{q},e_{a})$. The learning objective is formulated in a contrastive approach that maximizes the log-likelihood of correct triples in the training set, which amounts to minimizing:
$\mathcal{L}_{\text{train}}=-\!\!\!\!\!\!\!\!\!\sum_{(e_{q},r_{q},e_{a})\in\mathcal{F}_{\text{train}}}$ $\Big[\log\sigma\big(s(e_{q},r_{q},e_{a})\big)$ (2)
$\!+$ $\!\!\!\!\!\!\!\!\sum_{e^{\prime}_{n}\in\mathcal{N}_{(e_{q},r_{q},e_{a})}}\!\!\!\!\!\!\!\!\log\big(1-\sigma(s(e_{q},r_{q},e^{\prime}_{n}))\big)\Big],$
where $\sigma$ is a sigmoid function. GraphOracle retains this framework and introduces a Relation-Dependency Graph pre-training objective that endows $\phi$ with universal semantics, enabling zero-shot generalization to both unseen entities and unseen relation vocabularies.
4 The Proposed Method
In order to enable fully inductive KG reasoning and improve the generalization ability of models across KGs, the key is to generalize the dependencies among relations for different KGs. To achieve this goal, we firstly introduce Relational Dependency Graph (RDG), which explicitly models how relations depend on each other, in Section 4.1 Based on RDG, we propose a query-dependent multi-head attention mechanism to learn relation representations from a weighted combination of precedent relations in Section 4.2. Subsequently, in Section 4.3, we introduce the approach where entity representations are represented with the recursive function (1) in the original KGs by using the relation representations just obtained. The overview of our approach is shown in Fig 1.
4.1 RDG Construction
To build an effective KGFM capable of cross KG generalization, we must capture the fundamental dependent patterns through which one relation can be represented by a combination of others. Our key innovation is reparameterizing entity–relation interactions as a relation-dependency interaction structure that explicitly captures how relations influence each other.
Given a KG $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{F})$, we construct a RDG $\mathcal{G}^{\mathcal{R}}$ through a structural transformation. First, we define a relation adjacency operator $\Phi:\mathcal{F}\rightarrow\mathcal{R}\times\mathcal{R}$ that extracts transitive relation dependencies:
$\Phi(\mathcal{F})=\!\!\!\!\!\!\bigcup\limits_{e,e^{\prime},e^{\prime\prime}\in\mathcal{V}}\!\!\!\!\{(r_{i},r_{j})\mid(e,r_{i},e^{\prime})\in\mathcal{F}\land(e^{\prime},r_{j},e^{\prime\prime})\in\mathcal{F}\}.$
Then, the RDG is defined as $\mathcal{G}^{\mathcal{R}}=(\mathcal{R},\mathcal{E}^{\mathcal{R}})$, where the node set $\mathcal{R}$ contains all the relations and the edge sets $\mathcal{E}^{\mathcal{R}}=\Phi(\mathcal{F})$ includes relation dependencies induced by entity-mediated pathways. This transformation alters the conceptual framework, shifting from an entity-centric perspective to a relation-interaction manifold where compositional connections between relations become explicit. Each directed edge $(r_{i},r_{j})$ in $\mathcal{G}^{\mathcal{R}}$ represents a potential relation-dependency pathway, indicating that relation $r_{i}$ preconditions relation $r_{j}$ through their sequential interaction over a shared entity context. The edge structure encodes compositional relational semantics, capturing how one relation may influence the probability or applicability of another when they occur in sequence.
To incorporate the hierarchical and compositional nature of relation interactions, we define a partial ordering function $\tau:\mathcal{R}\rightarrow\mathbb{R}$ that assigns each relation a position in a relation precedence structure. This ordering is derived from the KG’s inherent structure through rigorous topological analysis of relation co-occurrence patterns and functional dependencies. Relations that serve as logical precursors in inference chains are assigned lower $\tau$ values, thereby establishing a directed acyclic structure in the relation graph that reflects the natural flow of information propagation. Using this ordering, we define the set of preceding relations for any relation $r_{v}$ as:
$\mathcal{N}^{\text{past}}(r_{v})\!=\!\{r_{u}\!\in\!\mathcal{R}\mid\!(r_{u},r_{v})\!\in\!\mathcal{E}^{\mathcal{R}}\text{ and }\tau(r_{u})\!<\!\tau(r_{v})\}.\!\!$ (3)
This formulation enables us to capture the directional dependency patterns where relations with lower positions in the hierarchy systematically precede and inform relations with higher $\tau$. By explicitly modeling these precedence relationships, our framework can identify and leverage compositional reasoning patterns that remain invariant across domains, enhancing the generalization capabilities.
4.2 Relation Representation Learning on RDG
Building on the constructed RDG $\mathcal{G}^{\mathcal{R}}$, we develop a representation mechanism that captures the contextualized semantics of relations conditioned on a specific query. Given a query relation $r_{q}$, we introduce an RDG aggregation mechanism to compute $d$-dimensional relation-node representations $\bm{R}_{q}\in\mathbb{R}^{|\mathcal{R}|\times d}$ conditional on $r_{q}$.
Following Eq. (1). we apply a labeling initialization to distinguish the query relation node $r_{q}$ in $\mathcal{G}^{\mathcal{R}}$. Then employ multi-head attention relation‑dependency message passing over the graph:
$\bm{h}_{r_{v}\mid r_{q}}^{0}=\texttt{INDICATOR}_{r}(r_{v},r_{q})=\delta_{r_{v},r_{q}}\cdot\bm{1}^{d},\quad r_{v}\in\mathcal{G}^{\mathcal{R}}$ (4)
$\bm{h}^{\ell}_{r_{v}\mid r_{q}}=\sigma\Big(\frac{1}{H}\sum_{h=1}^{H}\Big[\quad\;\bm{W}_{1}^{\ell,h}\!\!\!\!\sum_{r_{u}\in\mathcal{N}^{\text{past}}(r_{v})}\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{u}\mid r_{q}}$
$\quad\quad\quad\quad\quad+\bm{W}_{2}^{\ell,h}\,\hat{\alpha}_{r_{v}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{v}\mid r_{q}}\Big]\Big),$
where $\delta_{r_{v},r_{q}}=1$ if $v=q$, and 0 otherwise. $H$ is the number of attention heads, and $\bm{W}^{\ell,h}_{1},\bm{W}^{\ell,h}_{2}\in\mathbb{R}^{d\times d}$ are head-specific parameter matrices. The relation‑dependency attention weight $\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}$ captures the directional influence of relation $r_{u}$ on relation $r_{v}$, computed as:
$\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}=\frac{\exp\big(\bm{a}^{T}(\bm{W}^{h}_{\alpha}\bm{h}_{r_{u}}^{\ell-1}\|\bm{W}_{\alpha}^{h}\bm{h}_{r_{v}}^{\ell-1})\big)}{\sum_{r_{w}\in\mathcal{N}^{\text{past}}(r_{v})}\exp\big(\bm{a}^{T}(\bm{W}_{\alpha}^{h}\bm{h}_{r_{w}}^{\ell-1}\|\bm{W}_{\alpha}^{h}\bm{h}_{r_{v}}^{\ell-1})\big)},$
where $\bm{a}\in\mathbb{R}^{2d}$ is a learnable attention parameter vector, $\|$ denotes vector concatenation, and $\bm{W}^{h}_{\alpha}\in\mathbb{R}^{d\times d}$ are head-specific trainable projection matrix. The neighborhood function $\mathcal{N}^{\text{past}}(r_{v})$ enforces the relation‑dependency ordering of relations as defined in Eq. (3).
After $L_{r}$ layers of message passing, the final relation representation incorporates both local and higher-order dependencies $\bm{R}_{q}=\{\bm{h}_{r|r_{q}}^{L_{r}}\mid r\in\mathcal{R}\}$.
4.3 Entity Representation Learning on the Original KG
After obtaining the relation representations $\bm{R}_{q}$ from RDG conditioned on $r_{q}$, we obtain query-dependent entity representations by conducting message passing over the original KG structures. This approach enables effective reasoning across both seen and unseen entities and relations.
For a given query $(e_{q},r_{q},?)$, we compute entity representations recursively with Eq. (1) through the KG $\mathcal{G}$. The initial representations $\bm{h}^{0}_{e|q}=\bm{1}$ if $e=e_{q}$, and otherwise $\bm{0}$. At each layer $\ell$, the representation of an entity $e$ is computed as:
$\begin{array}[]{ll}\!\!\!\!\bm{h}^{\ell}_{e|q}=\delta\!\left(\bm{W}^{\ell}\!\cdot\!\sum_{(e_{s},r,e)\in\mathcal{F}_{\text{train}}}\!\alpha^{\ell}_{e_{s},r|q}\!\big(\bm{h}^{\ell-1}_{e_{s}|q}+\bm{h}^{L_{r}}_{r|r_{q}}\big)\!\right)\!,\!\!\end{array}$ (5)
where $\delta(\cdot)$ is a non-linear activation, and the attention weight $\alpha^{\ell}_{e_{s},r|q}$ is computed as:
$\alpha_{e_{s},r|q}^{\ell}=\sigma\Big((\bm{w}_{\alpha}^{\ell})^{\top}\text{ReLU}\big(\bm{W}_{\alpha}^{\ell}\cdot(\bm{h}^{\ell-1}_{e_{s}|q}\|\bm{h}^{L_{r}}_{r|r_{q}}\|\bm{h}^{L_{r}}_{r_{q}|r_{q}})\big)\Big).$
where $\bm{w}_{\alpha}^{\ell}\in\mathbb{R}^{d}$ and $\bm{W}_{\alpha}^{\ell}\in\mathbb{R}^{d\times 3d}$ are learnable parameters, $\sigma$ is the sigmoid function and $\cdot$ denotes the standard matrix-vector multiplication.
We iterate Eq. (5) for $L_{e}$ steps and use the final layer representation $\bm{h}_{e|q}^{L_{e}}$ for scoring each entity $e\in\mathcal{V}$. The critical idea here is replacing the learnable relation embeddings $\bm{r}$ with the contextualized relation embedding $\bm{h}^{L_{r}}_{r|r_{q}}$ from our RDG, enabling fully inductive reasoning (Time complexity of the GraphOracle model is given in Appendix C, and the Theoretical analysis on its expressiveness and generalization is given in Appendix I).
Figure 2: Comparison of the MRR performance (the larger the better) between GraphOracle and supervised SOTA methods across various datasets. Note that Amazon-book uses NDCG@20 due to its adaptation to the recommendation task.
4.4 Training Details
All the learnable parameters such as $\bigl\{\bm{W}^{h}_{O},\;\bm{W}^{\ell,h}_{1},\;\bm{W}^{\ell,h}_{2},\;\bm{W}^{h}_{\alpha},\;\bm{a},\;\bm{W}^{\ell},\;\bm{W}^{\ell}_{\alpha},\;\bm{w}^{\ell}_{\alpha},\;\bm{w}^{L}\bigr\}$ are trained end-to-end by minimizing the loss function Eq. (2). GraphOracle adopts a sequential multi‑dataset pre‑train $\rightarrow$ fine‑tune paradigm to acquire a general relation-dependency graph representation across KGs $\{\mathcal{G}_{1},\dots,\mathcal{G}_{K}\}$. For each graph $\mathcal{G}_{k}$, we optimize the regularized objective $\mathcal{L}^{(k)}=\mathcal{L}_{\text{task}}^{(k)}+\lambda_{k}\,\bigl\lVert\Theta\bigr\rVert_{2}^{\;2},$ where $\mathcal{L}_{\text{task}}^{(k)}$ denotes the task‑specific loss on $\mathcal{G}_{k}$ (e.g., Eq. (2)), $\Theta$ represents all learnable parameters, and $\lambda_{k}$ controls the strength of L2 regularization. Early stopping technique is used for each graph once validation MRR fails to improve for several epochs. This iterative pre‑train process, together with our relation‑dependency graph encoder, equips GraphOracle with strong cross‑domain generalization. When adapting GraphOracle to new KGs, we firstly build the RDG and then support two inference paradigms:
  • Zero-shot Inference. The pre-trained model is directly applied to unseen KGs without tuning.
  • Fine-tuning. For more challenging domains, we fine-tune the pre-trained parameters on the target KG $\mathcal{G}_{\text{target}}$ for a limited number of epochs $E_{\text{fine-tune}}\ll E_{\text{train}}$.
Model Transductive Entity Inductive Fully Inductive Cross-domain
Metric MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
Supervised SOTA 0.4185 0.4715 0.5771 0.4915 0.4730 0.6296 0.4593 0.2942 0.6246 0.2964 0.2149 0.4458
GraphOracle 0.4486 0.5550 0.6111 0.5449 0.5684 0.6722 0.5203 0.3688 0.7279 0.3759 0.2744 0.5485
Improvement 7.19% 17.70% 5.89% 10.86% 20.16% 6.77% 13.28% 25.36% 16.54% 26.82% 27.70% 23.03%
Table 2: Average performance comparison between GraphOracle and Supervised SOTA under four generalization settings.
Figure 3: Perturbation Analysis of RDG Edges by Attention-Derived Importance Scores.
Figure 4: Impact of Number of Pre-trained Datasets on Zero-Shot Evaluation Metrics.
5 Experiment
To evaluate the comprehensive capabilities of GraphOracle as a Foundation Model for KG reasoning, we formulate the following research questions: RQ1: How does GraphOracle model perform compared with state-of-the-art models on diverse KGs and cross-domain datasets? RQ2: How do different relation-dependency patterns contribute to GraphOracle’s performance? RQ3: To what extent can external information enhance the performance of GraphOracle? RQ4: How do the components and configurations contribute to the performance?
5.1 Experimental Setup
Datasets
We conduct comprehensive experiments on $60$ KGs, which we classify into three categories according to their properties (Details are given in Appendix D.):
  • Transductive and Inductive datasets. To ensure fair comparison, we follow the same dataset settings as ULTRA (galkin2024ultra), TRIX (zhang2025trix), and KG-ICL (cui2024promptkg), including 16 transductive, 18 entity-inductive, and 23 fully-inductive datasets, totaling 57 in all.
  • Cross domain datasets. (i) Biomedical Datasets: We use biomedical KG PrimeKG (chandak2023building) to examine the cross-domain capabilities of GraphOracle. We finetune with 80% samples in raw PrimeKG, and validate with 10% samples. When testing on the remaining 10%, we specially focus on the predictions for triplets: (Protein, Interacts_with, BP/MF/CC), (Drug, Indication, Disease), (Drug, Target, Protein), (Protein, Associated_with, Disease), (Drug, Contraindication, Disease). (ii) Recommendation domain: We transform the Amazon-book (wang2019kgat) dataset into a pure KG reasoning dataset to adapt to the KGs Reasoning field by defining the interactions between users and items as a new relation in the KG. (iii) Geographic datasets (GeoKG) (geonames) . (Detail process are given in Appendix F.)
Pretrain and Finetune.
GraphOracle is pre-trained on three general KGs (NELL-995, CoDEx-Medium, FB15k-237) to capture diverse relational structures and reasoning patterns. It takes 150,000 training steps with batch size of 32 using AdamW optimizer (loshchilov2019decoupled) on a single A6000 (48GB) GPU. For cross-domain adaptation, we employ a lightweight fine-tuning approach that updates only the final layer parameters while freezing the pre-trained representations. The finetune process only takes $1\sim 2$ epochs to achieve the best results. The pre-training process takes approximately 36 hours, while fine-tuning requires only $15\sim 60$ minutes depending on the target dataset. Detailed hyperparameters, architecture specifications, and training configurations are provided in Appendix E.
Baselines
We compare the proposed GraphOracle with (i) Transductive: ConvE (ConvE), QuatE (QuatE), DuASE (li2024duase) and BioBRIDGE (wang2024biobridge); (ii) Entity inductive: MINERVA (MINERVA), DRUM (DRUM), AnyBURL (meilicke2020reinforced), RNNLogic (RNNLogic), RLogic (RLogic) GraphRulRL (mai2025graphrulrl), CompGCN (CompGCN), NBFNet (zhu2021neural), RED-GNN (RED-GNN), A*Net (Zhu2023AStarNet), Adaprop (Adaprop) and one-shot-subgraph (one-shot-subgraph); (iii) Fully inductive: INGRAM (lee2023ingram), ULTRA (galkin2024ultra), TRIX (zhang2025trix) and KG-ICL (cui2024promptkg). The results for the baseline methods were either directly obtained from the original publications or reproduced using the official source code provided by the authors. Due to page limitations, some other baselines can be found in INGRAM (lee2023ingram), BioBRIDGE (wang2024biobridge) and KUCNet (liu2024kucnet).
5.2 Overall Performance (RQ1)
Figure 5: Comparison on PrimeKG: Evaluating GraphOracle Enhanced by External Entity information (GraphOracle+).
Models Nell-100 WK-100 FB-100 YAGO3-10 GeoKG
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
GraphOracle 0.702 0.623 0.905 0.417 0.192 0.711 0.576 0.407 0.812 0.696 0.672 0.807 0.639 0.552 0.793
W/o RDG 0.376 0.278 0.493 0.132 0.102 0.243 0.237 0.149 0.432 0.593 0.545 0.712 0.521 0.452 0.614
W/o multi-head 0.598 0.534 0.817 0.302 0.39 0.619 0.492 0.321 0.724 0.629 0.576 0.722 0.566 0.477 0.646
$\text{Graph}_{\text{INGRAM}}$ 0.478 0.375 0.663 0.189 0.092 0.389 0.398 0.283 0.557 0.478 0.512 0.625 0.485 0.471 0.613
$\text{Graph}_{\text{ULTRA}}$ 0.548 0.490 0.720 0.201 0.105 0.466 0.465 0.297 0.679 0.587 0.583 0.725 0.545 0.491 0.645
$\text{Message}_{\text{INGRAM}}$ 0.517 0.426 0.726 0.271 0.111 0.518 0.428 0.289 0.635 0.539 0.473 0.674 0.496 0.481 0.624
$\text{Message}_{\text{ULTRA}}$ 0.569 0.502 0.741 0.282 0.124 0.597 0.468 0.305 0.672 0.503 0.539 0.669 0.457 0.479 0.636
Table 3: Ablation Analysis of GraphOracle’s Core Architectural Components across Five Benchmark Datasets.
The main experimental results, presented in Fig. 2, illustrate the performance of GraphOracle following pre-training on three KGs and a brief fine-tuning phase of just two epochs across 60 distinct datasets (comprehensive results are available in Appendix G). A salient finding is that GraphOracle consistently outperforms the supervised SOTA across all evaluated baseline datasets and metrics, as detailed in Table 2. This robust performance underscores the overall efficacy of our methodology, with particularly notable improvements observed in the more challenging scenarios. The results demonstrate substantial gains across all reasoning types, with the most pronounced improvements occurring in fully inductive and cross-domain settings, where the model must generalize to entirely unseen entities or domains—scenarios that represent the most stringent tests of a model’s reasoning capabilities.
5.3 Relation-Dependency Pattern Analysis (RQ2)
To investigate whether GraphOracle truly internalizes the compositionality of the relations, that is, the way complex relations are constructed systematically from simpler ones, we performed a series of perturbation analyzes on the learned RDG. First, we calculated relation attention weights using Eq. (4.2), averaged over the WN18RR and YAGO3-10 datasets, to assign an importance score to each relation pair, reflecting its contribution to downstream reasoning. Subsequently, during inference, we systematically disabled specific subsets of edges based on these attention weights: (i) the top-5 and top-10 most important (highest attention) compositional relation pairs; (ii) the bottom-5 and bottom-10 least important (lowest attention) pairs; and (iii) 5 and 10 randomly selected pairs.
As shown in Fig. 4, removing the highly-ranked compositional edges—those encoding key multi-hop templates essential for composing higher-level relations—causes a sharp decline in MRR on both WN18RR and YAGO3-10. This confirms that GraphOracle heavily relies on these identified compositional pathways for its predictions. Conversely, suppressing a small number of low-importance edges sometimes leads to slight performance improvements, suggesting that these weaker compositional cues might act as semantic noise. Perturbations involving randomly removed edges result in only moderate performance degradation. This underscores the idea that it is not merely the quantity of relations but the specific, learned compositional interactions between them that are crucial for GraphOracle’s reasoning process. These findings collectively substantiate that GraphOracle’s predictions are rooted in the compositional structure of its RDG, rather than relying on isolated relation statistics.
5.4 Compatible with Additional Initial Information (RQ3)
To explore the potential of external information in enhancing KG reasoning, we introduced an improved entity initialization strategy. This involved incorporating modality-specific encoded features as initial entity vectors, moving beyond standard random initialization. The resulting model, denoted as GraphOracle+, leverages foundation model embeddings to create more semantically rich entity representations (details are provided in Appendix H). As demostrated in Fig 5 (details are given in Tabel 17), experimental evaluations on the PrimeKG dataset show that GraphOracle+ achieves consistent performance gains across all metrics. Notably, MRR scores improved by 15% for protein-biological process prediction and 14% for protein-molecular function prediction. These results affirm that GraphOracle’s framework significantly benefits from integrating external information. In an era increasingly influenced by large language models, the capacity for flexible incorporation of diverse information sources is crucial for advancing generalization and adaptability, especially within specialized and complex domains such as biomedicine.
5.5 Ablation Study (RQ4)
To rigorously evaluate the contribution of each architectural component within GraphOracle, we conducted an extensive series of ablation experiments. We first investigate the impact of removing the RDG and the effect of reducing the number of attention heads $H$ in Eq. (4) from eight to one. The results are detailed in Table 3. Clearly, eliminating the RDG or the multi‑head attention mechanism causes a marked decline in all evaluation metrics, highlighting their indispensibility to GraphOracle’s performance.
In addition, we quantified how the breadth of pre‑training data affects zero‑shot performance. Specifically, we pre‑trained on one to six heterogeneous datasets and evaluated the resulting checkpoints on unseen Table 3’s graphs. The averaged results, depicted in Fig. 4 (implementation details are reported in Appendix E), reveal that zero-shot performance saturates once three diverse datasets are included in the pre-training mixture. Incorporating additional datasets beyond this point yields no further significant gains. We conjecture that after this point the model has already encountered a sufficiently rich spectrum of relational patterns, and subsequent datasets may introduce largely redundant or potentially noisy signals.
Furthermore, to underscore the unique contributions of our proposed mechanisms, we compared GraphOracle’s relation graph construction and message passing techniques against those employed by INGRAM and ULTRA. For this, we created variants where $\text{Graph}_{\text{INGRAM}}$ denotes using INGRAM’s method for relation graph construction,and $\text{Message}_{\text{INGRAM}}$ signifies adopting INGRAM’s message passing scheme (similarly for ULTRA). As detailed in Table 3, substituting either GraphOracle’s graph construction or its message passing method with those from INGRAM or ULTRA resulted in a substantial reduction in performance. This comparative analysis further substantiates the effectiveness and integral role of each distinct component within the GraphOracle framework.
6 Conclusion
In this work, we introduced GraphOracle, a relation-centric foundation model for unifying reasoning across heterogeneous KGs. By converting KGs into RDG, our approach explicitly encodes compositional patterns among relations, yielding domain-invariant embeddings. Experiments on 60 diverse benchmarks showed consistent state-of-the-art performance, improving mean reciprocal rank by up to 16.8% over baselines with minimal adaptation. We also demonstrated that GraphOracle’s performance can be enhanced by integrating external information through GraphOracle+, which leverages foundation model embeddings for improved initialization. Ablation studies confirmed the essential contributions of the relation-dependency graph and multi-head attention components. These findings establish relation-dependency pre-training as a scalable approach toward universal KG reasoning, opening new avenues for cross-domain applications.
A Reproducibility and Code
To ensure reproducibility, we provide the complete source code of the GraphOracle framework in the supplementary materials.
B Limitations and Future Work
Despite GraphOracle’s strong performance across various KG reasoning tasks, several limitations merit acknowledgment. The computational complexity of relation-dependency graph construction scales with the number of relations, which may present challenges for KGs with extremely high relation cardinality, potentially degrading efficiency for graphs with millions of distinct relations. Additionally, our approach currently focuses primarily on the topological structure of relation interactions and may not fully leverage all semantic nuances present in complex domain-specific knowledge, despite partial mitigation through GraphOracle+’s incorporation of external embeddings. Furthermore, while GraphOracle demonstrates strong zero-shot and few-shot capabilities, its performance still benefits from fine-tuning on target domains, indicating that truly universal KG reasoning remains challenging, particularly for highly specialized domains with unique relation structures.
The current work opens several promising directions for future research. Extending GraphOracle to incorporate multimodal knowledge sources represents a compelling direction where future architectures could jointly reason over textual descriptions, visual attributes, and graph structure to create more comprehensive knowledge representations, particularly valuable in domains like biomedicine where protein structures, medical images, and text reports contain complementary information. Additionally, developing temporal extensions to GraphOracle that model relation-dependency dynamics and knowledge evolution patterns would enable reasoning about causality, trends, and temporal dependencies between relations, addressing the static nature of current KGs. As KGs grow to include millions of relations, future research could explore techniques for automatically discovering relation taxonomies and leveraging them to create more efficient and scalable message-passing architectures through hierarchical abstractions of relation-dependencies.
C Time Complexity Analysis
GraphOracle’s overall time complexity consists of two parts: a one‑time preprocessing cost and a per‑query inference cost. Building the Relation–Dependency Graph (RDG) by scanning all triples once requires $\mathcal{O}(|F|),$ where $|F|$ is the total number of triples. For each query $(e_{q},r_{q},?)$, the $L_{R}$ layers of relation‑level message passing on the RDG incur $\mathcal{O}(L_{R}\,|E_{R}|\,d),$ where $|E_{R}|$ is the number of edges in the RDG and $d$ is the hidden dimension; in practice $|E_{R}|\ll|F|$ and $L_{R}\leq 3$, so this cost is small. Subsequently, the $L_{E}$ layers of entity‑level message passing propagate representations across an average branching factor $b$, costing $\mathcal{O}(L_{E}\,b\,d).$ With typical settings $L_{E}\leq 3$ and $b<30$, this yields sub‑millisecond latency per query. Hence the end‑to‑end per‑query complexity is $\mathcal{O}\bigl(L_{R}\,|E_{R}|\,d+L_{E}\,b\,d\bigr),$ while preprocessing remains $\mathcal{O}(|F|)$. Thanks to small constant depths and modest branching, GraphOracle achieves near‑linear scalability and memory‑efficient inference on large, heterogeneous knowledge graphs.
D Statistics of Datasets
We choose the same datasets as ULTRA (galkin2024ultra), TRIX (zhang2025trix) and KG-ICL (cui2024promptkg). Details of the used Knowledge Graph datasets are given in Table 4 and Table 5. Furthermore, we first create three cross-domain datasets for cross-domain Knowledge Graph Reasoning, whose details can be found in Appendix F.
Table 4: Statistics of the KG datasets. $Q_{\text{tra}}$, $Q_{\text{val}}$, $Q_{\text{tst}}$ are the query triplets used for reasoning.
Dataset Reference # Entity # Relation $|\mathcal{E}|$ $|Q_{\text{tra}}|$ $|Q_{\text{val}}|$ $|Q_{\text{tst}}|$ Supervised SOTA
WN18RR Dettmers et al. (2017) (dettmers2017convolutional) 40.9k 11 65.1k 21.7k 3.0k 3.1k one-shot-subgraph (one-shot-subgraph)
FB15k237 Toutanova and Chen (2015) (toutanova2015observed) 14.5k 237 204.1k 68.0k 17.5k 20.4k Adaprop (Adaprop)
NELL-995 Xiong et al.(2017) (xiong2017deeppath) 74.5k 200 112.2k 37.4k 543 2.8k Adaprop (Adaprop)
YAGO3-10 Suchanek et al.(2007) (suchanek2007yago) 123.1k 37 809.2k 269.7k 5.0k 5.0k one-shot-subgraph (one-shot-subgraph)
Nell-V1 Teru et al. (2020) (TER2020) 3.1k 14 5.5k 4.7k 0.4k 0.4k KG-ICL (cui2024promptkg)
Nell-V2 Teru et al. (2020) (TER2020) 2.6k 88 10.1k 8.2k 90.9k 1.0k KG-ICL (cui2024promptkg)
Nell-V3 Teru et al. (2020) (TER2020) 4.6k 142 20.1k 16.4k 1.9k 1.9k KG-ICL (cui2024promptkg)
Nell-V4 Teru et al. (2020) (TER2020) 2.1k 76 9.3k 7.5k 0.9k 0.9k KG-ICL (cui2024promptkg)
WN-V1 Teru et al. (2020) (TER2020) 2.7k 9 6.7k 5.4k 0.6k 0.6k KG-ICL (cui2024promptkg)
WN-V2 Teru et al. (2020) (TER2020) 7.0k 10 20.0k 15.3k 1.8k 1.9k KG-ICL (cui2024promptkg)
WN-V3 Teru et al. (2020) (TER2020) 12.1k 11 32.2k 25.9k 3.1k 3.2k KG-ICL (cui2024promptkg)
WN-V4 Teru et al. (2020) (TER2020) 3.9k 9 9.8k 7.9k 0.9k 1.0k KG-ICL (cui2024promptkg)
FB-V1 Teru et al. (2020) (TER2020) 1.6k 180 5.3k 4.2k 0.5k 0.5k KG-ICL (cui2024promptkg)
FB-V2 Teru et al. (2020) (TER2020) 2.6k 200 12.1k 9.7k 1.2k 1.2k KG-ICL (cui2024promptkg)
FB-V3 Teru et al. (2020) (TER2020) 3.7k 215 22.4k 18.0k 2.2k 2.2k KG-ICL (cui2024promptkg)
FB-V4 Teru et al. (2020) (TER2020) 4.7k 219 33.9k 27.2k 3.4k 3.4k KG-ICL (cui2024promptkg)
Nell-25 Lee et al. (2023) (lee2023ingram) 5.2k 146 19.1k 17.6k 0.7k 0.7k KG-ICL (cui2024promptkg)
Nell-50 Lee et al. (2023) (lee2023ingram) 5.3k 150 19.3k 17.6k 0.9k 0.9k KG-ICL (cui2024promptkg)
Nell-75 Lee et al. (2023) (lee2023ingram) 3.3k 138 12.3k 11.1k 6.1k 6.1k KG-ICL (cui2024promptkg)
Nell-100 Lee et al. (2023) (lee2023ingram) 2.1k 99 9.4k 7.8k 0.8k 0.8k KG-ICL (cui2024promptkg)
WK-25 Lee et al. (2023) (lee2023ingram) 13.9k 67 44.1k 41.9k 1.1k 1.1k KG-ICL (cui2024promptkg)
WK-50 Lee et al. (2023) (lee2023ingram) 16.3k 102 88.9k 82.5k 3.2k 3.2k KG-ICL (cui2024promptkg)
WK-75 Lee et al. (2023) (lee2023ingram) 8.1k 77 31.0k 28.7k 1.1k 1.1k KG-ICL (cui2024promptkg)
WK-100 Lee et al. (2023) (lee2023ingram) 15.9k 103 58.9k 49.9k 4.5k 4.5k KG-ICL (cui2024promptkg)
FB-25 Lee et al. (2023) (lee2023ingram) 8.7k 233 103.0k 91.6k 5.7k 5.7k KG-ICL (cui2024promptkg)
FB-50 Lee et al. (2023) (lee2023ingram) 8.6k 228 93.1k 85.4k 3.9k 3.9k KG-ICL (cui2024promptkg)
FB-75 Lee et al. (2023) (lee2023ingram) 6.9k 213 69.0k 62.8k 3.1k 3.1k KG-ICL (cui2024promptkg)
FB-100 Lee et al. (2023) (lee2023ingram) 6.5k 202 67.5k 62.8k 2.3k 2.3k KG-ICL (cui2024promptkg)
PrimeKG Chandak et al. (2023) (chandak2023building) 85.0k 14 3911.9k 3129.8k 391.2k 391.2k Adaprop (Adaprop)
Amazon-book Wang et al. (2019) (wang2019kgat) 3404.2k 40 3404.2k 3210.3k 98.0k 95.9k KUCNet (liu2024kucnet)
GeoKG GeoNames Team (2025) (geonames) 2054.2k 681 2784.5k 2673.1k 55.7k 55.7k one-shot-subgraph (one-shot-subgraph)
Table 5: Statistics of the KG datasets. $Q_{\text{tra}}$, $Q_{\text{val}}$, $Q_{\text{tst}}$ are the query triplets used for reasoning.
Dataset Reference # Entity # Relation $|Q_{\text{tra}}|$ $|Q_{\text{val}}|$ $|Q_{\text{tst}}|$ Supervised SOTA
CoDEx Small Dettmers et al. (2020) 2.0k 42 32.9k 1.8k 1.8k ULTRA (galkin2024ultra)
CoDEx Medium Dettmers et al. (2020) 17.1k 51 185.6k 10.3k 10.3k KG-ICL (cui2024promptkg)
CoDEx Large Dettmers et al. (2020) 78.0k 69 551.2k 30.6k 30.6k KG-ICL (cui2024promptkg)
WDsinger Lv et al. (2020) 10.3k 135 16.1k 2.2k 2.2k TRIX (zhang2025trix)
NELL23k Lv et al. (2020) 22.9k 200 25.4k 5.0k 5.0k KG-ICL (cui2024promptkg)
FB15k237_10 Lv et al. (2020) 11.5k 237 27.2k 15.6k 18.2k KG-ICL (cui2024promptkg)
FB15k237_20 Lv et al. (2020) 13.2k 237 54.4k 17.0k 20.0k KG-ICL (cui2024promptkg)
FB15k237_50 Lv et al. (2020) 14.1k 237 136.1k 17.4k 20.3k ULTRA (galkin2024ultra)
DBpedia100k Ding et al. (2018) 99.6k 470 597.6k 50.0k 50.0k TRIX (zhang2025trix)
AristoV4 Chen et al. (2021) 45.0k 1605 242.6k 20.0k 20.0k TRIX (zhang2025trix)
ConceptNet100k Malaviya et al. (2020) 78.3k 34 100.0k 1.2k 1.2k KG-ICL (cui2024promptkg)
Hetionet Himmelstein et al. (2017) 45.2k 24 2025.2k 112.5k 112.5k ULTRA (galkin2024ultra)
E Comprehensive Training Details
Sequential multi‑dataset schedule.
Given $K$ KGs $\{\mathcal{G}_{1},\ldots,\mathcal{G}_{K}\}$ sorted by domain diversity, we train GraphOracle sequentially from $\mathcal{G}_{1}$ to $\mathcal{G}_{K}$. Parameters are rolled over between datasets to accumulate relational knowledge.
Dataset‑specific hyper‑parameters.
Each dataset $\mathcal{G}_{k}$ employs a tuple $\Theta_{k}=\{\alpha_{k},\lambda_{k},\gamma_{k},d_{k}^{(\mathrm{h})},d_{k}^{(\mathrm{a})},\delta_{k},\mathcal{A}_{k},L_{k}\}$ denoting learning rate, $\ell_{2}$ regularization, decay factor, hidden dimension, attention dimension, dropout rate, activation function, and layer count, respectively. Values are chosen via grid search on the validation split of $\mathcal{G}_{k}$.
Objective.
$\mathcal{L}^{(k)}$ $=\mathcal{L}_{\mathrm{train}}^{(k)}+\lambda_{k}\lVert\Theta\rVert_{2}^{2},$ (6)
$\mathcal{L}_{\mathrm{train}}^{(k)}$ $=\mathbb{E}_{(h,r,t)\sim\mathcal{D}_{k}}\bigl[-\log p(t\mid h,r;\Theta)\bigr],$ (7)
with negative sampling ratio $N_{\text{neg}}=64$.
Learning‑rate decay and early stopping.
At epoch $\epsilon$ we apply $\alpha_{k}^{(t+1)}=\gamma_{k}\!\cdot\!\alpha_{k}^{(t)}$; training terminates when validation MRR has not improved for $10$ epochs.
Parameter transfer.
After convergence on $\mathcal{G}_{k}$, we initialize the next run via
$\Theta_{k+1}^{(0)}=\mathcal{T}(\Theta_{k}^{\star}),$ (8)
where $\mathcal{T}$ preserves (i) relation‑dependency graph weights and (ii) shared layer norms, while re‑initializing dataset‑specific embeddings.
Table 6: Graphs in different pre-training mixtures.
Dataset 1 2 3 4 5 6
WN18RR
CoDEx-Medium
FB15k237
NELL995
PrimeKG
GeoKG
Batch size 50 10 10 5 2 2
Table 7: Hyperparameters used across different datasets.
Datasets Learning Rate Act Entity Layer Relation Layer Hidden Dim Batch Size
WN18RR 0.003 idd 5 3 64 50
FB15k237 0.0009 relu 4 4 48 10
NELL-995 0.0011 relu 5 4 48 5
YAGO3-10 0.001 relu 7 4 64 5
NELL-100 0.0016 relu 5 3 48 10
NELL-75 0.0013 relu 5 3 48 10
NELL-50 0.0015 tanh 5 3 48 10
NELL-25 0.0016 relu 5 3 48 10
WK-100 0.0027 relu 5 3 48 10
WK-75 0.0018 relu 5 3 48 10
WK-50 0.0022 relu 5 3 48 10
WK-25 0.0023 idd 5 3 48 10
FB-100 0.0043 relu 5 3 48 10
FB-75 0.0037 relu 5 3 48 10
FB-50 0.0008 relu 5 3 48 10
FB-25 0.0005 tanh 5 3 16 24
WN-V1 0.005 idd 5 3 64 100
WN-V2 0.0016 relu 5 4 48 20
WN-V3 0.0014 tanh 5 4 64 20
WN-V4 0.006 relu 5 3 32 10
FB-V1 0.0092 relu 5 3 32 20
FB-V2 0.0077 relu 3 3 48 10
FB-V3 0.0006 relu 3 3 48 20
FB-V4 0.0052 idd 5 4 48 20
NL-V1 0.0021 relu 5 3 48 10
NL-V2 0.0075 relu 3 3 48 100
NL-V3 0.0008 relu 3 3 16 10
NL-V4 0.0005 tanh 5 4 16 20
PrimeKG 0.00016 relu 5 4 16 2
Amazon-book 0.0002 idd 5 3 32 3
GeoKG 0.0005 relu 4 3 16 2
We perform a grid search and use the Optuna library to search for the optimal hyperparameters. Table 7 presents the choices of hyperparameters on all datasets, and Table 6 presents the choices of datasets in the pre-training process.
F Cross-domain Dataset Processing Details
The datasets used in this paper are all open source and can be obtained from:
F.1 Processing Details of Biomedical Domain
Table 8: Comparison on PrimeKG. Best performance is highlighted with bold, and the second best is underlined. GraphOracle-S means the GraphOracle was trained from scratch, while GraphOracle-F means the GraphOracle was trained by fine-tuning.
[…]
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
Embedding TransE 0.034 0.023 0.052 0.046 0.034 0.073 0.044 0.027 0.074 0.017 0.010 0.030 0.033 0.022 0.053 0.024 0.014 0.045 0.010 0.005 0.019
TransR 0.045 0.030 0.068 0.060 0.044 0.095 0.048 0.030 0.081 0.053 0.032 0.093 0.069 0.046 0.112 0.028 0.016 0.052 0.029 0.015 0.055
TransSH 0.044 0.029 0.067 0.061 0.045 0.096 0.057 0.035 0.096 0.026 0.016 0.046 0.043 0.028 0.070 0.024 0.014 0.045 0.014 0.007 0.026
TransD 0.043 0.029 0.065 0.059 0.044 0.093 0.053 0.033 0.090 0.022 0.013 0.039 0.049 0.032 0.079 0.024 0.014 0.045 0.013 0.007 0.025
ComplEx 0.084 0.056 0.128 0.100 0.074 0.158 0.099 0.061 0.167 0.042 0.025 0.074 0.079 0.052 0.128 0.059 0.034 0.110 0.048 0.025 0.091
DistMult 0.054 0.036 0.082 0.089 0.066 0.141 0.095 0.059 0.161 0.025 0.015 0.044 0.044 0.029 0.071 0.033 0.019 0.062 0.047 0.025 0.089
RotatE 0.079 0.053 0.120 0.119 0.088 0.188 0.107 0.066 0.181 0.150 0.090 0.264 0.125 0.083 0.203 0.070 0.041 0.131 0.076 0.040 0.144
BioBRIDGE 0.136 0.091 0.207 0.326 0.241 0.515 0.319 0.198 0.539 0.189 0.113 0.333 0.172 0.114 0.279 0.084 0.049 0.157 0.081 0.043 0.153
GNNs NBFNet 0.279 0.187 0.424 0.335 0.248 0.529 0.321 0.199 0.543 0.169 0.101 0.297 0.156 0.103 0.253 0.200 0.116 0.374 0.139 0.074 0.263
RED-GNN 0.284 0.190 0.432 0.341 0.252 0.539 0.327 0.203 0.553 0.172 0.103 0.303 0.159 0.105 0.258 0.203 0.118 0.379 0.142 0.075 0.268
A*Net 0.317 0.212 0.482 0.381 0.282 0.602 0.365 0.226 0.617 0.192 0.115 0.338 0.177 0.117 0.287 0.227 0.132 0.424 0.158 0.084 0.299
AdaProp 0.334 0.224 0.508 0.402 0.297 0.635 0.385 0.239 0.651 0.202 0.121 0.356 0.187 0.124 0.303 0.239 0.139 0.447 0.167 0.089 0.316
one-shot-subgraph 0.231 0.155 0.351 0.278 0.206 0.439 0.266 0.165 0.450 0.140 0.084 0.246 0.129 0.085 0.209 0.165 0.096 0.308 0.115 0.061 0.217
INGRAM 0.269 0.180 0.409 0.324 0.240 0.512 0.310 0.192 0.524 0.163 0.098 0.287 0.151 0.100 0.245 0.193 0.112 0.361 0.134 0.071 0.253
ULTRA 0.313 0.210 0.476 0.376 0.278 0.594 0.360 0.223 0.608 0.189 0.113 0.333 0.175 0.116 0.284 0.224 0.130 0.419 0.156 0.083 0.295
GraphOracle-S 0.392 0.251 0.500 0.423 0.314 0.698 0.431 0.263 0.692 0.223 0.132 0.397 0.197 0.139 0.334 0.262 0.154 0.487 0.176 0.103 0.345
GraphOracle-F 0.498 0.323 0.684 0.499 0.366 0.748 0.475 0.325 0.752 0.268 0.175 0.448 0.232 0.167 0.378 0.299 0.187 0.531 0.192 0.135 0.380
Dataset Creation
Table 9: Statistical analysis of node and edge distribution in the original PrimeKG dataset and our processed KG used for training. “Original” represents the raw PrimeKG data; “Processed” indicates our filtered KG; “Dropped” shows the number of entities removed during preprocessing.
Type Modality Original Processed Dropped Percent dropped
Nodes biological process 28,642 27,478 1,164 4.06%
protein 27,671 19,162 8,509 30.75%
disease 17,080 17,080 0 0.00%
molecular function 11,169 10,966 203 1.82%
drug 7,957 6,948 1,009 12.68%
cellular component 4,176 4,013 163 3.90%
Summation 96,695 85,647 11,048 11.43%
Type Relation Original Processed Dropped Percent Dropped
Edges drug_drug 2,672,628 2,241,466 431,162 16.13%
protein_protein 642,150 629,208 12,942 2.02%
bioprocess_protein 289,610 272,642 16,968 5.86%
cellcomp_protein 166,804 149,504 17,300 10.37%
disease_protein 160,822 155,924 4,898 3.05%
molfunc_protein 139,060 133,522 5,538 3.98%
bioprocess_bioprocess 105,772 99,630 6,142 5.81%
disease_disease 64,388 64,388 0 0.00%
contraindication 61,350 60,130 1,220 1.99%
drug_protein 51,306 47,614 3,692 7.20%
molfunc_molfunc 27,148 26,436 712 2.62%
indication 18,776 17,578 1,198 6.38%
cellcomp_cellcomp 9,690 9,200 490 5.06%
off-label use 5,136 4,998 138 2.69%
Summation 4,414,640 3,912,240 502,400 11.38%
Our research establishes a framework for integrating uni-modal foundation models through KG simplification. As show in Table LABEL:tab:detail_of_BioKG, for experimental efficiency, we refined the KG to include six key modalities. These retained modalities—protein, disease, drug, and gene ontology terms—represent the core biomedical entities crucial for addressing real-world applications including drug discovery, repurposing, protein-protein interaction analysis, protein function prediction, and drug-target interaction modeling. Table 9 presents a comprehensive comparison between the original and our processed KG.
For classic KG reasoning dataset, we retain only the IDs of entities and relations, while preserving their modalities as auxiliary information. These modalities are used solely for categorizing relation types during statistical analysis and are not involved in the message passing process. The dataset is partitioned into training, validation, and test sets with a standard split of 80%, 10%, and 10%, respectively.
External Information-Enrichment Dataset Creation
The initial PrimeKG dataset aggregates biomedical entities from numerous sources. To enhance the utility of this dataset for our purposes, we enriched entities with essential properties and established connections to external knowledge bases, removing entities lacking required attributes.
Protein Entities
The original PrimeKG contains 27,671 protein entries. We implemented a mapping procedure to associate these proteins with UniProtKB/Swiss-Prot sequence database through the UniProt ID mapping service (https://www.uniprot.org/id-mapping). This procedure yielded 27,478 protein sequences successfully matched with gene identifiers.
Further analysis of the unmapped entries revealed that most corresponded to non-protein-coding genetic elements (including pseudogenes, rRNA, and ncRNA genes), which do not produce functional proteins. Given our focus on protein-centric applications, excluding these entries was appropriate.
Drug Entities
From the initial 7,957 drug entries in PrimeKG, we performed identity matching against the DrugBank database (https://go.drugbank.com/drugs). During this process, we removed drugs lacking SMILES structural notation, resulting in 6,948 validated drug entities for our training dataset.
Gene Ontology Terms
The biological process, molecular function, and cellular component categories comprise the Gene Ontology (GO) terminology in our dataset. We utilized AmiGO to extract detailed definitions of these GO terms through their identifiers (https://amigo.geneontology.org/amigo/search/ontology). This process allowed us to incorporate 27,478 biological process terms, 10,966 molecular function terms, and 4,013 cellular component terms into our training dataset.
Disease Entities
Disease descriptions were directly adopted from the PrimeKG dataset, allowing us to retain all 17,080 disease entities without modification for training purposes.
F.2 Processing Detail of Recommendation Domain
Table 10: Comparison of the performance of different methods on Amazon-book. The best performance is marked in bold and the second best performance is underlined. The GraphOracle-S means the GraphOracle was train from scratch while GraphOracle-F means the GraphOracle was trained by finetune.
Method MF FM NFM RippleNet KGNN-LS CKAN KGIN CKE R-GCN KGAT PPR PathSim RED-GNN KUCNet GraphOracle-S GraphOracle-F
Recall@20 0 0.0026 0.0006 0.0011 0.0001 0.0005 0.0868 0 0.0001 0.0001 0.0301 0.2053 0.2187 0.2237 0.2453 0.3142
NDCG@20 0 0.0010 0.0003 0.0005 0.0001 0.0003 0.0446 0 0.0001 0.0001 0.0167 0.1491 0.1633 0.1685 0.1987 0.2591
Input: $\mathcal{R}$ (relation list), $\mathcal{T}_{\text{train}}$ (training triples), $\mathcal{K}_{G}$ (original KG), $\mathcal{T}_{\text{test}}$ (test triples)
Output: $\mathcal{R}^{\prime}$ (extended relations), $\mathcal{K}^{\prime}_{G}$ (enriched KG), $\mathcal{M}_{E}$ (entity–ID map)
1
2 /* Phase I: Relation Ontology Extension */;
3 $\bm{\Omega}\leftarrow\textsc{ReadRelations}(\mathcal{R})$;
4 $r_{\max}\leftarrow\max_{r\in\bm{\Omega}}\textsc{id}(r)$;
5 $\text{id}_{\text{purchase}}\leftarrow r_{\max}+1$;
6 $\mathcal{R}^{\prime}\leftarrow\bm{\Omega}\cup\{(\text{``purchase''},\text{id}_{\text{purchase}})\}$;
7 PersistRelationSet()($\mathcal{R}^{\prime}$);
8
9 /* Phase II: Semantic Triple Generation */;
10 $\mathcal{P}\leftarrow\emptyset$;
11 foreach $\sigma\in\mathcal{T}_{\text{train}}$ do
12 $\mathcal{E}\leftarrow\textsc{TokenizeEntities}(\sigma)$;
 $u\leftarrow\mathcal{E}[0]$;
  // user
 $\mathcal{I}_{u}\leftarrow\mathcal{E}[1:]$;
  // items
13 foreach $i\in\mathcal{I}_{u}$ do
14    $\mathcal{P}\leftarrow\mathcal{P}\cup\{(u,\text{id}_{\text{purchase}},i)\}$;
15
16   end foreach
17
18 end foreach
19
20 /* Phase III: KG Augmentation */;
21 $\mathcal{K}_{\text{ori}}\leftarrow\textsc{ExtractTriples}(\mathcal{K}_{G})$;
22 $\mathcal{K}^{\prime}_{G}\leftarrow\mathcal{K}_{\text{ori}}\cup\mathcal{P}$;
[…]
24
25 /* Phase IV: Entity Canonicalization */;
26 $\mathcal{E}_{\text{train}}\leftarrow\{s,o\mid(s,\_,o)\in\mathcal{K}^{\prime}_{G}\}$;
27 $\mathcal{E}_{\text{test}}\leftarrow\bigcup_{\sigma\in\mathcal{T}_{\text{test}}}\textsc{TokenizeEntities}(\sigma)$;
28 $\mathcal{E}_{\text{uni}}\leftarrow\mathcal{E}_{\text{train}}\cup\mathcal{E}_{\text{test}}$;
29 $\mathcal{E}_{\text{ord}}\leftarrow\textsc{TopologicalSort}(\mathcal{E}_{\text{uni}})$;
30 for $j\leftarrow 0$ to $|\mathcal{E}_{\text{ord}}|-1$ do
31 $\Phi(\mathcal{E}_{\text{ord}}[j])\leftarrow j$;
32
33 end for
[…]
return $\mathcal{R}^{\prime},\mathcal{K}^{\prime}_{G},\mathcal{M}_{E}$
Algorithm 1 KG Enrichment & Entity Canonicalization
Theoretical Framework and Implementation Principles
The algorithm presented herein delineates a comprehensive methodology for KG enrichment and entity canonicalization within the recommendation domain. This approach operates through a multi-phase framework that transforms heterogeneous data sources into a unified semantic representation amenable to graph-based recommendation algorithms.
Phase I (Relation Ontology Extension) introduces a formal extension of the relational schema $\mathcal{R}$ with a domain-specific ”purchase” relation. Let $\bm{\Omega}=\{(r_{1},id_{1}),(r_{2},id_{2}),...,(r_{n},id_{n})\}$ represent the initial relation set. The algorithm derives $r_{max}=\max_{r\in\bm{\Omega}}(\text{id}(r))$ and establishes $\text{id}_{purchase}=r_{max}+1$, thus creating an extended relation ontology $\mathcal{R}^{\prime}=\bm{\Omega}\cup\{(\text{"purchase"},\text{id}_{purchase})\}$. This expansion facilitates the semantic representation of user-item interactions within the KG structure.
Phase II (Semantic Triple Generation) transforms implicit user-item interactions into explicit RDF-compatible triples. For each user $u\in\mathcal{U}$ and their associated items $\mathcal{I}_{u}\subseteq\mathcal{I}$, where $\mathcal{U}$ and $\mathcal{I}$ denote the user and item entity spaces respectively, the algorithm constructs a set of purchase triples defined by:
$\mathcal{P}=\bigcup_{u\in\mathcal{U}}\bigcup_{i\in\mathcal{I}_{u}}\{(u,\text{id}_{purchase},i)\}$ (9)
These triples codify the user-item engagement patterns within the formalism of a KG, enabling the integration of collaborative filtering signals with semantic relationships.
Phase III (KG Augmentation) implements the fusion of the original KG $\mathcal{K}_{original}$ with the newly derived purchase triples $\mathcal{P}$. The enriched KG $\mathcal{K}^{\prime}_{G}$ is formalized as:
$\mathcal{K}^{\prime}_{G}=\mathcal{K}_{original}\cup\mathcal{P}$ (10)
This augmentation creates a multi-relational graph structure that encapsulates both semantic domain knowledge and behavioral interaction patterns.
Phase IV (Entity Canonicalization) establishes a unified reference framework for all entities across both training and evaluation datasets. The algorithm constructs the universal entity set $\mathcal{E}_{universal}=\mathcal{E}_{train}\cup\mathcal{E}_{test}$, where $\mathcal{E}_{train}$ comprises entities appearing in $\mathcal{K}^{\prime}_{G}$ and $\mathcal{E}_{test}$ consists of entities present in the test dataset. A bijective mapping function $\Phi:\mathcal{E}_{universal}\rightarrow\{0,1,...,|\mathcal{E}_{universal}|-1\}$ is implemented to assign canonical integer identifiers to each entity, facilitating efficient indexing and dimensional reduction.
The canonicalization process ensures consistent entity representation across both KG construction and recommendation evaluation, mitigating potential entity alignment issues and optimizing computational efficiency. The resulting entity mapping $\mathcal{M}_{E}=\{(e,\Phi(e))|e\in\mathcal{E}_{universal}\}$ enables seamless integration of the KG with neural recommendation architectures that typically require numerical entity representations.
This algorithmic framework yields a semantically enriched KG with standardized entity references, establishing the foundation for knowledge-aware recommendation algorithms that can simultaneously leverage collaborative signals and semantic relationships to generate contextualized and interpretable recommendations.
F.3 Processing Detail of Geographic Datasets (GeoKG)
Table 11: Comparison of the performance of different methods on GeoKG. The best performance is marked in bold and the second best performance is underlined. The GraphOracle-S means the GraphOracle was train from scratch while GraphOracle-F means the GraphOracle was trained by finetune.
Method TransE TransR TransSH TransD ComplEx DistMult RotatE NBFNet RED-GNN ANet AdaProp one-shot-subgraph ULTRA GraphOracle-S GraphOracle-F
MRR 0.013 0.032 0.019 0.021 0.075 0.084 0.093 0.425 0.486 0.473 0.493 0.473 0.528 0.539 0.639
H@1 0.028 0.064 0.039 0.042 0.088 0.097 0.104 0.403 0.443 0.428 0.450 0.428 0.486 0.493 0.552
H@10 0.046 0.088 0.048 0.053 0.103 0.125 0.176 0.534 0.566 0.564 0.573 0.586 0.627 0.654 0.793
Input: KG $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$, prune ratio $\rho$, visibility ratio $\theta$, split ratios $\bm{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3})$, weight $w$
Output: Splits $\{\mathcal{G}_{i}\}_{i=1}^{3}$, entity map $\pi_{\mathcal{E}}$, relation map $\pi_{\mathcal{R}}$
1
2 /* Phase I: Relation-Aware Pruning */;
3 $\mathcal{T}\leftarrow\textsc{Unique}(\mathcal{T})$;
4 Group $\mathcal{T}$ by relation: $\mathcal{T}_{r}$;
5 Compute normalized degree $\bar{\bm{d}}(v)=\bm{d}(v)\big/\max_{u}\bm{d}(u)$;
6 $\mathcal{T}^{\rho}\leftarrow\emptyset$;
7 foreach $r\in\mathcal{R}$ do
8 foreach $(h,r,t)\in\mathcal{T}_{r}$ do
9    $\Upsilon(h,r,t)\leftarrow w\cdot\frac{\bar{\bm{d}}(h)+\bar{\bm{d}}(t)}{2}$;
10
11   end foreach
12 $\mathcal{T}^{\rho}\leftarrow\mathcal{T}^{\rho}\cup\textsc{Top}_{\rho}(\mathcal{T}_{r},\Upsilon)$;
13
14 end foreach
15Define pruned KG $\mathcal{G}^{\rho}$ from $(\mathcal{E}^{\rho},\mathcal{R}^{\rho},\mathcal{T}^{\rho})$;
16
17 /* Phase II: Visibility Partitioning */;
18 Randomly split $\mathcal{E}^{\rho}$ and $\mathcal{R}^{\rho}$ into seen/unseen by $\theta$;
19 $\mathcal{T}_{\text{train}}\leftarrow$ triples whose $h,t,r$ are all seen;
20 $\mathcal{T}_{\text{eval}}\leftarrow\mathcal{T}^{\rho}\setminus\mathcal{T}_{\text{train}}$;
21
22 /* Phase III: Distribution Enforcement */;
23 Target $|\mathcal{T}_{\text{train}}|=\alpha_{1}|\mathcal{T}^{\rho}|$;
24 Balance $\mathcal{T}_{\text{train}}$ and $\mathcal{T}_{\text{eval}}$ via random moves;
25 Split $\mathcal{T}_{\text{eval}}$ into validation/test by $(\alpha_{2},\alpha_{3})$, ensuring disjointness;
26
27 /* Phase IV: Finalization */;
28 Form $\mathcal{G}_{1}=(\mathcal{E}^{\text{train}},\mathcal{R}^{\text{train}},\mathcal{T}_{\text{train}})$,   $\mathcal{G}_{2}=(\mathcal{E}^{\text{valid}},\mathcal{R}^{\text{valid}},\mathcal{T}_{\text{valid}})$,   $\mathcal{G}_{3}=(\mathcal{E}^{\text{test}},\mathcal{R}^{\text{test}},\mathcal{T}_{\text{test}})$;
29 Build index maps $\pi_{\mathcal{E}},\pi_{\mathcal{R}}$ by ascending order of IDs;
return $\{\mathcal{G}_{i}\}_{i=1}^{3},\;\pi_{\mathcal{E}},\;\pi_{\mathcal{R}}$
Algorithm 2 Relation-Balanced Pruning & Partitioning for Geographic KG
Fully-Inductive Geographic KG Dataset Construction.
We propose a comprehensive framework for constructing, pruning, and partitioning geographic KGs, designed to ensure relation balance, semantic diversity, and full inductiveness while preserving critical structural information. Our framework systematically addresses five major challenges: (i) maintaining relation diversity, (ii) preserving structural integrity, (iii) controlling visibility of entities and relations, (iv) enforcing strict data partitioning, and (v) ensuring data integrity via rigorous duplicate prevention.
The core innovation lies in the relation-balanced pruning strategy introduced in Phase I. Instead of applying a global importance metric across all triples, we stratify the pruning process by relation type. For each relation $r\in\mathcal{R}$, we select the top $\rho=7.5\%$ most important triples based on a relation-specific scoring function $\Upsilon_{r}$, which estimates the structural importance of triples involving entities $h$ and $t$ via:
$\Upsilon_{r}(h,r,t)=w\cdot\frac{\bar{\bm{d}}(h)+\bar{\bm{d}}(t)}{2}$ (11)
where $\bar{\bm{d}}(\cdot)$ denotes the normalized degree of an entity. This ensures that the pruned graph $\mathcal{G}^{\rho}$ maintains balanced semantic representation across both frequent and rare relations, avoiding dominance by high-frequency edges.
To guarantee inductiveness, Phase II performs visibility-controlled partitioning by randomly assigning 70% of entities and relations to the training set. All training triples are composed solely of these “seen” elements, while validation and test triples each include at least one “unseen” entity or relation. This design ensures a fully-inductive setup, where no inference triple shares entities or relations with the training set, thereby enabling robust assessment of generalization to entirely new graph components.
Phase III enforces the 80%-10%-10% train-validation-test ratio by adjusting assignments from Phase II when necessary, while strictly ensuring that each triple appears in only one split. Phase IV finalizes the dataset by constructing the three graph partitions and generating sequential ID mappings for all entities and relations.
Overall, our framework yields a semantically diverse, structurally meaningful, and fully-inductive geographic KG dataset that is reduced to 7.5% of the original size. The relation-aware pruning and controlled visibility mechanisms work in tandem to ensure both data compactness and inductive generalization capability, while robust duplicate handling preserves data integrity throughout the pipeline.
G Complete Experimental Results
In this section, we report the comprehensive experimental results of our study. Table 12 presents the performance of GraphOracle on transductive benchmarks, while Table 13, Table 14 and Table 15 summarize the results on entity-inductive and fully-inductive settings, respectively. Cross-domain evaluations are provided in Table 8, Table 10, and Table 11. Furthermore, Table 17 illustrates the enhanced performance of GraphOracle+ when external information is incorporated. Across all datasets and evaluation scenarios, GraphOracle consistently outperforms existing baselines by a notable margin, highlighting the robustness and effectiveness of our proposed approach.
Table 12: Comparison of GraphOracle with other reasoning methods in the transductive setting. Best performance is indicated by the bold face numbers, and the underline means the second best. “–” means unavailable results
Type Model WN18RR FB15k237 NELL-995 YAGO3-10
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
Non-GNN ConvE 0.427 39.2 49.8 0.325 23.7 50.1 0.511 44.6 61.9 0.520 45.0 66.0
QuatE 0.480 44.0 55.1 0.350 25.6 53.8 0.533 46.6 64.3 0.379 30.1 53.4
RotatE 0.477 42.8 57.1 0.337 24.1 53.3 0.508 44.8 60.8 0.495 40.2 67.0
MINERVA 0.448 41.3 51.3 0.293 21.7 45.6 0.513 41.3 63.7
DRUM 0.486 42.5 58.6 0.343 25.5 51.6 0.532 46.0 66.2 0.531 45.3 67.6
AnyBURL 0.471 44.1 55.2 0.301 20.9 47.3 0.398 27.6 45.4 0.542 47.7 67.3
RNNLogic 0.483 44.6 55.8 0.344 25.2 53.0 0.416 36.3 47.8 0.554 50.9 62.2
RLogic 0.477 44.3 53.7 0.310 20.3 50.1 0.416 25.2 50.4 0.360 25.2 50.4
DuASE 0.489 44.8 56.9 0.329 23.5 51.9 0.423 37.2 59.2 0.473 38.7 62.8
GraphRulRL 0.483 44.6 54.1 0.385 31.4 57.5 0.425 27.8 52.7 0.432 35.4 51.7
GNNs CompGCN 0.479 44.3 54.6 0.355 26.4 53.5 0.463 38.3 59.6 0.421 39.2 57.7
NBFNet 0.551 49.7 66.6 0.415 32.1 59.9 0.525 45.1 63.9 0.550 47.9 68.6
RED-GNN 0.533 48.5 62.4 0.374 28.3 55.8 0.543 47.6 65.1 0.559 48.3 68.9
A*Net 0.549 49.5 65.9 0.411 32.1 58.6 0.549 48.6 65.2 0.563 49.8 68.6
AdaProp 0.562 49.9 67.1 0.417 33.1 58.5 0.554 49.3 65.5 0.573 51.0 68.5
ULTRA 0.480 47.9 61.4 0.368 33.9 56.4 0.509 46.2 66.0 0.557 53.1 71.0
one-shot-subgraph 0.567 51.4 66.6 0.304 22.3 45.4 0.547 48.5 65.1 0.606 54.0 72.1
TRIX 0.514 48.1 61.1 0.366 32.5 55.9 0.506 44.2 64.8 0.541 47.3 70.2
KG-ICL 0.536 49.6 63.7 0.376 32.7 53.8 0.534 46.7 67.2 0.545 47.4 68.8
GraphOracle 0.675 61.7 76.2 0.471 39.6 66.4 0.621 56.3 75.1 0.696 67.2 80.7
Table 13: Comparison of GraphOracle with other reasoning methods in the entity inductive setting. Best performance is indicated by the bold face numbers, and the underline means the second best.
WN18RR FB15k-237 NELL-995
Models V1 V2 V3 V4 V1 V2 V3 V4 V1 V2 V3 V4
MRR RuleN 0.668 0.645 0.368 0.624 0.363 0.433 0.439 0.429 0.615 0.385 0.381 0.333
Neural LP 0.649 0.635 0.361 0.628 0.325 0.389 0.400 0.396 0.610 0.361 0.367 0.261
DRUM 0.666 0.646 0.380 0.627 0.333 0.395 0.402 0.410 0.628 0.365 0.375 0.273
GraIL 0.627 0.625 0.323 0.553 0.279 0.276 0.251 0.227 0.481 0.297 0.322 0.262
CoMPILE 0.577 0.578 0.308 0.548 0.287 0.276 0.262 0.213 0.330 0.248 0.319 0.229
NBFNet 0.684 0.652 0.425 0.604 0.307 0.369 0.331 0.305 0.584 0.410 0.425 0.287
RED-GNN 0.701 0.690 0.427 0.651 0.369 0.469 0.445 0.442 0.637 0.419 0.436 0.363
AdaProp 0.733 0.715 0.474 0.662 0.310 0.471 0.471 0.454 0.644 0.452 0.435 0.366
ULTRA 0.685 0.679 0.411 0.614 0.509 0.524 0.504 0.496 0.757 0.575 0.563 0.469
TRIX 0.705 0.682 0.425 0.650 0.515 0.525 0.501 0.493 0.804 0.571 0.571 0.551
KG-ICL 0.762 0.721 0.503 0.683 0.531 0.568 0.537 0.525 0.841 0.641 0.631 0.594
GraphOracle 0.807 0.793 0.569 0.762 0.619 0.631 0.694 0.658 0.864 0.684 0.659 0.619
Hit@1 (%) RuleN 63.5 61.1 34.7 59.2 30.9 34.7 34.5 33.8 54.5 30.4 30.3 24.8
Neural LP 59.2 57.5 30.4 58.3 24.3 28.6 30.9 28.9 50.0 24.9 26.7 13.7
DRUM 61.3 59.5 33.0 58.6 24.7 28.4 30.8 30.9 50.0 27.1 26.2 16.3
GraIL 55.4 54.2 27.8 44.3 20.5 20.2 16.5 14.3 42.5 19.9 22.4 15.3
CoMPILE 47.3 48.5 25.8 47.3 20.8 17.8 16.6 13.4 10.5 15.6 22.6 15.9
NBFNet 59.2 57.5 30.4 57.4 19.0 22.9 20.6 18.5 50.0 27.1 26.2 23.3
RED-GNN 65.3 63.3 36.8 60.6 30.2 38.1 35.1 34.0 52.5 31.9 34.5 25.9
AdaProp 66.8 64.2 39.6 61.1 19.1 37.2 37.7 35.3 52.2 34.4 33.7 24.7
ULTRA 61.5 58.7 33.5 58.7 32.2 39.9 40.5 37.2 50.7 35.8 36.4 28.8
TRIX 63.9 58.4 34.7 59.3 32.9 39.8 40.7 37.0 53.9 35.5 36.9 31.7
KG-ICL 65.4 61.7 36.9 60.5 41.1 43.8 42.6 39.6 59.4 39.6 41.2 35.6
GraphOracle 76.8 79.8 47.3 67.8 40.4 51.9 52.8 50.9 65.6 52.6 51.3 44.9
Hit@10 (%) RuleN 73.0 69.4 40.7 68.1 44.6 59.9 60.0 60.5 76.0 51.4 53.1 48.4
Neural LP 77.2 74.9 47.6 70.6 46.8 58.6 57.1 59.3 87.1 56.4 57.6 53.9
DRUM 77.7 74.7 47.7 70.2 47.4 59.5 57.1 59.3 87.3 54.0 57.7 53.1
GraIL 76.0 77.6 40.9 68.7 42.9 42.4 42.4 38.9 56.5 49.6 51.8 50.6
CoMPILE 74.7 74.3 40.6 67.0 43.9 45.7 44.9 35.8 57.5 44.6 51.5 42.1
NBFNet 82.7 79.9 56.3 70.2 51.7 63.9 58.8 55.9 79.5 63.5 60.6 59.1
RED-GNN 79.9 78.0 52.4 72.1 48.3 62.9 60.3 62.1 86.6 60.1 59.4 55.6
AdaProp 86.6 83.6 62.6 75.5 55.1 65.9 63.7 63.8 88.6 65.2 61.8 60.7
ULTRA 79.3 77.9 54.6 72.0 67.0 71.0 66.3 68.4 87.8 76.1 75.5 73.3
TRIX 79.8 78.0 54.3 72.2 68.2 73.0 69.9 68.7 89.9 76.4 75.9 77.2
KG-ICL 82.7 78.7 62.6 74.9 70.0 76.8 70.4 70.6 99.5 83.5 79.9 80.2
GraphOracle 92.3 92.6 69.5 81.6 76.7 79.2 75.7 78.5 97.2 86.9 86.2 82.4
Table 14: Comparison of GraphOracle with other reasoning methods in fully-inductive setting. Best performance is indicated by the bold face numbers, and the underline means the second best. H@1” and H@10” are short for Hit@1 and Hit@10 (in percentage), respectively. –” means unavailable results.
Model Nell-100 Nell-75 Nell-50 Nell-25
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
GraIL 0.135 0.114 0.173 0.096 0.056 0.205 0.162 0.104 0.288 0.216 0.160 0.366
CoMPILE 0.123 0.071 0.209 0.178 0.093 0.361 0.194 0.125 0.330 0.189 0.115 0.324
SNRI 0.042 0.029 0.064 0.088 0.040 0.177 0.130 0.095 0.187 0.190 0.140 0.270
INDIGO 0.160 0.109 0.247 0.121 0.098 0.156 0.167 0.134 0.217 0.166 0.134 0.206
RMPI 0.220 0.136 0.376 0.138 0.061 0.275 0.185 0.109 0.307 0.213 0.130 0.329
CompGCN 0.008 0.001 0.014 0.014 0.003 0.025 0.003 0.000 0.005 0.006 0.000 0.010
NodePiece 0.012 0.004 0.018 0.042 0.020 0.081 0.037 0.013 0.079 0.098 0.057 0.166
NeuralLP 0.084 0.035 0.181 0.117 0.048 0.273 0.101 0.064 0.190 0.148 0.101 0.271
DRUM 0.076 0.044 0.138 0.152 0.072 0.313 0.107 0.070 0.193 0.161 0.119 0.264
BLP 0.019 0.006 0.037 0.051 0.012 0.120 0.041 0.011 0.093 0.049 0.024 0.095
QBLP 0.004 0.000 0.003 0.040 0.007 0.095 0.048 0.020 0.097 0.073 0.027 0.151
NBFNet 0.096 0.032 0.199 0.137 0.077 0.255 0.225 0.161 0.346 0.283 0.224 0.417
RED-GNN 0.212 0.114 0.385 0.203 0.129 0.353 0.179 0.115 0.280 0.214 0.166 0.266
RAILD 0.018 0.005 0.037
INGRAM 0.309 0.212 0.506 0.261 0.167 0.464 0.281 0.193 0.453 0.334 0.241 0.501
ULTRA 0.458 0.423 0.684 0.374 0.369 0.570 0.418 0.256 0.595 0.407 0.278 0.596
TRIX 0.482 0.437 0.691 0.351 0.325 0.525 0.405 0.213 0.555 0.377 0.262 0.589
KG-ICL 0.557 0.459 0.766 0.446 0.378 0.681 0.528 0.274 0.708 0.540 0.301 0.730
GraphOracle 0.702 0.623 0.905 0.612 0.423 0.923 0.589 0.421 0.868 0.579 0.389 0.923
Model WK-100 WK-75 WK-50 WK-25
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
CompGCN 0.003 0.000 0.009 0.015 0.003 0.028 0.003 0.001 0.002 0.009 0.000 0.020
NodePiece 0.007 0.002 0.018 0.021 0.003 0.052 0.008 0.002 0.013 0.053 0.019 0.122
NeuralLP 0.009 0.005 0.016 0.020 0.004 0.054 0.025 0.007 0.054 0.068 0.046 0.104
DRUM 0.010 0.004 0.019 0.020 0.007 0.043 0.017 0.002 0.046 0.064 0.035 0.116
BLP 0.012 0.003 0.025 0.043 0.016 0.089 0.041 0.013 0.092 0.125 0.055 0.283
QBLP 0.012 0.003 0.025 0.044 0.016 0.091 0.035 0.011 0.080 0.116 0.042 0.294
NBFNet 0.014 0.005 0.026 0.072 0.028 0.172 0.062 0.036 0.105 0.154 0.092 0.301
RED-GNN 0.096 0.070 0.136 0.172 0.110 0.290 0.058 0.033 0.093 0.170 0.111 0.263
RAILD 0.026 0.010 0.052
INGRAM 0.107 0.072 0.169 0.247 0.179 0.362 0.068 0.034 0.135 0.186 0.124 0.309
ULTRA 0.168 0.089 0.286 0.380 0.278 0.635 0.140 0.076 0.280 0.321 0.388 0.535
TRIX 0.188 0.093 0.290 0.368 0.254 0.513 0.166 0.078 0.313 0.300 0.354 0.401
KG-ICL 0.270 0.127 0.415 0.466 0.313 0.626 0.277 0.091 0.432 0.425 0.434 0.628
GraphOracle 0.417 0.192 0.711 0.469 0.345 0.698 0.362 0.101 0.498 0.582 0.460 0.780
Table 15: Comparison of GraphOracle with other reasoning methods in fully-inductive setting. Best performance is indicated by the bold face numbers, and the underline means the second best. H@1” and H@10” are short for Hit@1 and Hit@10 (in percentage), respectively. –” means unavailable results.
Model FB-100 FB-75 FB-50 FB-25
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
CompGCN 0.015 0.008 0.025 0.013 0.000 0.026 0.004 0.002 0.006 0.003 0.000 0.004
NodePiece 0.006 0.001 0.009 0.016 0.007 0.029 0.021 0.006 0.048 0.044 0.011 0.114
NeuralLP 0.026 0.007 0.057 0.056 0.030 0.099 0.088 0.043 0.184 0.164 0.098 0.309
DRUM 0.034 0.011 0.077 0.065 0.034 0.121 0.101 0.061 0.191 0.175 0.109 0.320
BLP 0.017 0.004 0.035 0.047 0.024 0.085 0.078 0.037 0.156 0.107 0.053 0.212
QBLP 0.013 0.003 0.026 0.041 0.017 0.084 0.071 0.030 0.147 0.104 0.043 0.226
NBFNet 0.072 0.026 0.154 0.089 0.048 0.166 0.130 0.071 0.259 0.224 0.137 0.410
RED-GNN 0.121 0.053 0.263 0.107 0.057 0.201 0.129 0.072 0.251 0.145 0.077 0.284
RAILD 0.031 0.016 0.048
INGRAM 0.223 0.146 0.371 0.189 0.119 0.325 0.117 0.067 0.218 0.133 0.067 0.271
ULTRA 0.444 0.287 0.643 0.400 0.269 0.598 0.334 0.275 0.538 0.383 0.242 0.635
TRIX 0.436 0.269 0.633 0.401 0.263 0.611 0.334 0.277 0.547 0.393 0.256 0.650
KG-ICL 0.499 0.307 0.719 0.458 0.274 0.664 0.384 0.291 0.598 0.434 0.279 0.694
GraphOracle 0.576 0.407 0.812 0.538 0.356 0.872 0.585 0.434 0.913 0.562 0.370 0.930
Table 16: Performance comparison among ULTRA, TRIX, KG-ICL, and GraphOracle across different datasets. Best results are in bold and second best are underlined.
Type Model ULTRA TRIX KG-ICL GraphOracle
MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10
Transductive CoDEx Small 0.490 0.686 0.484 0.676 0.479 0.662 0.512 0.697
CoDEx Medium 0.372 0.525 0.365 0.521 0.402 0.565 0.417 0.574
CoDEx Large 0.343 0.478 0.388 0.481 0.388 0.508 0.396 0.523
WDsinger 0.417 0.526 0.502 0.620 0.493 0.599 0.512 0.654
NELL23k 0.268 0.450 0.306 0.536 0.329 0.552 0.333 0.572
FB15k237_10 0.254 0.411 0.253 0.408 0.260 0.416 0.269 0.435
FB15k237_20 0.274 0.445 0.273 0.441 0.284 0.456 0.297 0.482
FB15k237_50 0.325 0.528 0.322 0.522 0.324 0.499 0.336 0.541
DBpedia100k 0.436 0.603 0.457 0.619 0.455 0.604 0.479 0.643
AristoV4 0.343 0.496 0.345 0.499 0.313 0.480 0.374 0.524
ConceptNet100k 0.310 0.529 0.340 0.564 0.371 0.584 0.386 0.602
Entity Inductive Hetionet 0.399 0.538 0.394 0.534 0.269 0.402 0.417 0.556
ILPC Small 0.303 0.453 0.310 0.455 0.316 0.473 0.339 0.497
ILPC Large 0.308 0.431 0.310 0.431 0.295 0.411 0.345 0.451
HM 1k 0.042 0.100 0.072 0.128 0.089 0.144 0.097 0.178
HM 3k 0.030 0.090 0.069 0.118 0.081 0.129 0.089 0.143
HM 5k 0.025 0.068 0.074 0.118 0.070 0.108 0.096 0.145
IndigoBM 0.432 0.639 0.436 0.645 0.440 0.641 0.483 0.697
Fully Inductive MT1 tax 0.330 0.459 0.397 0.508 0.411 0.521 0.491 0.568
MT1 health 0.380 0.467 0.376 0.457 0.387 0.479 0.405 0.501
MT2 org 0.104 0.170 0.098 0.162 0.100 0.171 0.132 0.193
MT2 sci 0.311 0.451 0.331 0.526 0.303 0.396 0.337 0.574
MT3 art 0.306 0.473 0.289 0.461 0.306 0.460 0.315 0.481
MT3 infra 0.657 0.807 0.672 0.810 0.676 0.808 0.697 0.829
MT4 sci 0.303 0.478 0.305 0.482 0.307 0.473 0.321 0.496
MT4 health 0.704 0.785 0.702 0.785 0.710 0.776 0.721 0.796
Metafam 0.997 1.000 0.997 1.000 1.000 1.000 1.000 1.000
FBNELL 0.481 0.661 0.478 0.655 0.516 0.699 0.523 0.732
NL-0 0.329 0.551 0.385 0.549 0.555 0.765 0.566 0.777
H Detail Design of GraphOracle+
Table 17: Performance Comparison on PrimeKG: Evaluating GraphOracle Enhanced by External Entity Initialization (GraphOracle+)
[…]
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
GraphOracle .498 .323 .684 .499 .366 .748 .475 .325 .752 .268 .175 .448 .232 .167 .378 .299 .187 .531 .192 .135 .380
GraphOracle+ .573 .371 .787 .574 .421 .860 .546 .374 .865 .308 .201 .515 .267 .192 .435 .344 .215 .611 .221 .155 .437
In real-world scenarios, users often query models with procedural or temporal “How”-type questions rather than isolated factual prompts. Script-based evaluation frameworks (li2025scedit) emphasize the importance of integrating external knowledge to support dynamic, multi-step reasoning. Motivated by this, we extend GraphOracle by incorporating modality-specific external features to enhance its representational capacity.
For each entity $e_{i}$, we update its information as $e_{i}=\{x^{i},c^{i}\}$ which combines an additional feature vector $x^{i}$ and a modality tag $c^{i}$. For example, in PrimeKG, a drug may be defined as $c^{i}=\text{"drug"}$ and $x^{i}=\text{"functional description of the drug"}$. To further enhance GraphOracle, we introduce a methodological extension by integrating external information, resulting in an improved variant termed GraphOracle+. Specifically, we incorporate multiple unimodal foundation models (uni-FMs), and by leveraging the embeddings generated by these uni-FMs, we effectively enrich the representations of individual nodes. In the current era of large-scale models, the ability to seamlessly integrate heterogeneous sources of information is of paramount importance. As described, for any two entities $e_{i}$ and $e_{j}$ originating from distinct modalities, we utilize modality-specific foundation models to encode their features. The initial embedding for entity $e_{i}$ under a query context $(e_{q},r_{q})$ is formulated as:
$\bm{h}_{e_{i}}(e_{q},r_{q})=\psi(x^{i},c^{i})$ (12)
where $\psi(x^{i},c^{i})$ denotes a modality-specific encoder selected based on the entity type $c^{i}$, and $x^{i}$ represents the raw input features of $e_{i}$. To unify embeddings produced by different unimodal encoders into a common representation space, we introduce a modality-aware projection function $\mathcal{T}(c^{i})$, which aligns each modality to a shared latent space. $\bm{h}_{e_{i}}^{0}(e_{q},r_{q})=\mathcal{T}(\bm{h}_{e_{i}}(e_{q},r_{q}),c^{i})\in\mathbb{R}^{d},$ where $c^{i}$ denotes the modality type of entity $e_{i}$, and $\mathcal{T}(\cdot,\cdot)$ ensures that all modality-specific outputs are projected into a unified $d$-dimensional space. Furthermore, to make the model modally-aware, we encode its modality type $c^{i}$ to obtain its modality embedding $\bm{c}^{i}$. The complete encoding process is:
$\begin{array}[]{ll}&\bm{h}^{0}_{e}(e_{q},r_{q},{c^{i}})=\mathcal{T}(\bm{h}_{e_{i}}(e_{q},r_{q}),c^{i})=\mathcal{T}(\psi(x^{i},c^{i}),c^{i}),\\[5.0pt] &\bm{h}^{\ell}_{e_{o}}(e_{q},r_{q},{c^{i}}^{\ell})=\delta\left(\bm{W}^{\ell}\cdot\sum_{(e_{s},r,e_{o})\in\hat{\mathcal{E}}^{\ell}_{e_{q}}}\alpha^{\ell}_{e_{s},r,e_{o}|r_{q}}\left(\bm{h}^{\ell-1}_{e_{s}}(e_{q},r_{q},{c^{i}}^{\ell-1})+\Psi({\bm{c}^{i}}^{\ell-1},{\bm{c}^{i}}^{\ell},\bm{h}^{\ell}_{r})\right)\right),\end{array}$
where ${\alpha}^{\ell}_{e_{s},r,e_{s}|r_{q}}$ is defined the same as Eq. (4.3), ${\bm{c}^{i}}^{\ell-1}$ and $\bm{c}^{i}$$^{\ell}$ is the modality embedding of nodes at $\ell-1$ and $\ell$ respectively, and $\Psi$ is a vanilla six-layer transformer model for bridging different modalities. Table 17 demonstrates the powerful performance of GraphOracle+ and proves the scalability of our model.
I Theoretical Analysis of the GraphOracle Model
In this section, we provide rigorous theoretical guarantees for the GraphOracle framework, analyzing its expressiveness, generalization capabilities, convergence properties, stability under perturbations, and relation‑dependency correctness.
I.1 Expressiveness and Representation Capacity
Theorem 1 (Representation Capacity).
The RDG representation in GraphOracle with L message passing layers can distinguish between any two non-isomorphic relation subgraphs with a maximum path length of L.
Proof.
We prove this by induction on the number of message passing layers L.
Base case ($\ell$=1): For $\ell=1$, the representation of relation r after one message passing layer is:
$\bm{h}^{\ell}_{r_{v}\mid r_{q}}=\sigma\left(\frac{1}{H}\sum_{h=1}^{H}\left[\bm{W}_{1}^{\ell,h}\sum_{r_{u}\in\mathcal{N}^{\text{past}}(r_{v})}\hat{\alpha}_{r_{u}r}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{u}\mid r_{q}}+\bm{W}_{2}^{\ell,h}\,\hat{\alpha}_{r_{v}r}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{v}\mid r_{q}}\right]\right),$
Since the initial representation $\bm{h}_{r|r_{q}}^{0}=\delta_{r,r_{q}}\cdot\bm{1}^{d}$ distinguishes the query relation from all others, and the attention weights $\hat{{\alpha}}_{r_{u}r}^{h}$ are distinct for different neighborhood configurations, non-isomorphic relation subgraphs of depth 1 will have distinct representations.
Inductive step: Assume the statement holds for $L=k$. For $L=k+1$, each relation’s representation now incorporates information from relations that are $k+1$ steps away. If two relation subgraphs are non-isomorphic within $k+1$ steps, either:
  1. 1.
    They were already non-isomorphic within $k$ steps, which by our inductive hypothesis leads to different representations, or The difference occurs exactly at step $k+1$, which will result in different inputs to the $(k+1)$-th layer message passing function, thus producing different representations.
Therefore, the theorem holds for all $L$. ∎
Theorem 2 (Expressive Power).
For any continuous function $f:\mathcal{X}\to\mathcal{Y}$ on compact sets $\mathcal{X}$ and $\mathcal{Y}$, there exists a GraphOracle model with sufficient width and depth that can approximate $f$ with arbitrary precision.
Proof.
The proof leverages the universal approximation theorem for neural networks. Our model consists of three components:
1. The RDG representation module:
$\bm{h}^{\ell}_{r_{v}\mid r_{q}}=\sigma\left(\frac{1}{H}\sum_{h=1}^{H}\left[\bm{W}_{1}^{\ell,h}\sum_{r_{u}\in\mathcal{N}^{\text{past}}(r_{v})}\hat{\alpha}_{r_{u}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{u}\mid r_{q}}+\bm{W}_{2}^{\ell,h}\,\hat{\alpha}_{r_{v}r_{v}}^{\ell,h}\,\bm{h}^{\ell-1}_{r_{v}\mid r_{q}}\right]\right),$
2. The universal entity representation module:
$\bm{h}^{\ell}_{e|q}=\delta\left(\bm{W}^{\ell}\cdot\sum_{(e_{s},r,e)\in\mathcal{F}_{\text{train}}}\alpha^{\ell}_{e_{s},r|r_{q}}\left(\bm{h}^{\ell-1}_{e_{s}|q}+\bm{h}^{L_{r}}_{r|r_{q}}\right)\right),$ (16)
3. The scoring function:
$s(e_{q},r_{q},e_{a})=\bm{w}_{s}^{\top}\bm{h}_{r_{q}}^{L}(e_{q},e_{a})$ (17)
Each component is constructed from differentiable functions that can be approximated by neural networks with sufficient capacity. By the universal approximation theorem, for any continuous function $f$ and any $\epsilon>0$, there exists a neural network that approximates $f$ within an error bound of $\epsilon$.
Therefore, with sufficient width (embedding dimension $d$) and depth (number of layers $L$), GraphOracle can approximate any continuous function over the KG with arbitrary precision. ∎
I.2 Generalization Bounds
Theorem 3 (Generalization Error Bound).
For a GraphOracle model with parameters $\Theta$ trained on a dataset $\mathcal{D}$ with $N$ triples sampled from a KG with $|\mathcal{V}|$ entities and $|\mathcal{R}|$ relations, the expected generalization error is bounded by:
$\mathbb{E}[\mathcal{L}_{test}(\Theta)-\mathcal{L}_{train}(\Theta)]\leq\mathcal{O}\left(\sqrt{\frac{\log(|\mathcal{V}|\cdot|\mathcal{R}|)}{N}}\right)$ (18)
Proof.
Let $\mathcal{H}$ be the hypothesis class of all possible GraphOracle models with fixed architecture. The VC-dimension of $\mathcal{H}$ can be bounded by $\mathcal{O}(p\log p)$, where $p$ is the number of parameters in the model, which is proportional to $|\mathcal{R}|\cdot d^{2}\cdot L$.
By standard results from statistical learning theory, the generalization error is bounded by:
$\mathbb{E}[\mathcal{L}_{test}(\Theta)-\mathcal{L}_{train}(\Theta)]\leq\mathcal{O}\left(\sqrt{\frac{VC(\mathcal{H})}{N}}\right)$ (19)
Substituting our bound on the VC-dimension:
$\mathbb{E}[\mathcal{L}_{test}(\Theta)-\mathcal{L}_{train}(\Theta)]\leq\mathcal{O}\left(\sqrt{\frac{|\mathcal{R}|\cdot d^{2}\cdot L\cdot\log(|\mathcal{R}|\cdot d^{2}\cdot L)}{N}}\right)$
Since $d$ and $L$ are fixed hyperparameters of the model, and $|\mathcal{R}|$ is bounded by the KG size, we can simplify this to:
$\mathbb{E}[\mathcal{L}_{test}(\Theta)-\mathcal{L}_{train}(\Theta)]\leq\mathcal{O}\left(\sqrt{\frac{\log(|\mathcal{V}|\cdot|\mathcal{R}|)}{N}}\right)$ (21)
This completes the proof. ∎
Theorem 4 (Inductive Generalization).
Let $\mathcal{G}_{train}=(\mathcal{V}_{train},\mathcal{R}_{train},\mathcal{F}_{train})$ and $\mathcal{G}_{test}=(\mathcal{V}_{test},\mathcal{R}_{test},\mathcal{F}_{test})$ be training and testing KGs. If the relation structures are similar, i.e., $d_{TV}(\mathcal{G}^{\mathcal{R}}_{train},\mathcal{G}^{\mathcal{R}}_{test})\leq\epsilon$, then the generalization error is bounded by:
$\mathcal{L}_{test}(\Theta)-\mathcal{L}_{train}(\Theta)\leq\mathcal{O}(\epsilon+\sqrt{\frac{\log(|\mathcal{V}_{train}|\cdot|\mathcal{R}_{train}|)}{N}})$ (22)
where $d_{TV}$ is the total variation distance between the relation graphs.
Proof.
We decompose the generalization error into two components:
$\mathcal{L}_{test}(\Theta)-\mathcal{L}_{train}(\Theta)=[\mathcal{L}_{test}(\Theta)-\mathcal{L}^{*}_{test}(\Theta)]+[\mathcal{L}^{*}_{test}(\Theta)-\mathcal{L}_{train}(\Theta)]$
where $\mathcal{L}^{*}_{test}(\Theta)$ is the expected loss under the optimal parameter setting for the test graph.
The first term represents the approximation error due to structural differences between train and test graphs, which is bounded by $\mathcal{O}(\epsilon)$ based on the similarity assumption.
The second term is the standard generalization error from the previous theorem.
Combining these bounds completes the proof. ∎
Figure 6: Comparison of the effects of building relationship graphs using different methods
J Comparison of GraphOracle with Supervised SOTA Methods
Table 18: The number of edges in the relation graph constructed by INGRAM, ULTRA, and GraphOracle.
Dataset # Relation INGRAM ULTRA GraphOracle
NL-25 146 1610 2300 797
NL-50 150 1748 2526 861
NL-75 138 1626 2336 787
NL-100 99 892 1159 416
WK-25 67 598 947 256
WK-50 102 1130 2164 508
WK-75 77 732 1253 313
WK-100 103 1052 1695 460
FB-25 233 7172 10479 3501
FB-50 228 6294 9300 3135
FB-75 213 5042 7375 2524
FB-100 202 4058 5728 2017
WN_V1 9 48 40 37
WN_V2 10 76 76 55
WN_V3 11 94 85 68
WN_V4 9 70 61 54
FB_V1 180 1622 2416 712
FB_V2 200 2692 4050 1237
FB_V3 215 3398 5015 1640
FB_V4 219 4624 7036 2231
NL_V1 14 122 170 51
NL_V2 88 1574 2065 842
NL_V3 142 1942 2558 1017
NL_V4 76 1296 1657 744
(a) KG1
(b) INGRAM
(c) ULTRA
(d) GraphOracle
(e) KG2
(f) INGRAM
(g) ULTRA
(h) GraphOracle
Figure 7: Comparison of GraphOracle’s graph construction with other methods
(a) Knowledge Graph
(b) Relation Graph
Figure 8: Illustration of GraphOracle’s Relation-Dependency Graph Construction
Fig. 7 illustrates the comparative transfer learning performance (KG1 $\rightarrow$ KG2) of GraphOracle, INGRAM, and ULTRA [^1] Fig. 8 shows the visualization process of GraphOracle extracting RDG.. Unlike INGRAM, which connects every pair of relations sharing an entity–thus producing an undirected $\mathcal{O}(|\mathcal{R}|^{2})$ co-occurrence graph with indiscriminate propagation that ignores directional dependencies–and ULTRA, which subdivides those links into four fixed head/tail interaction patterns (head-to-head, head-to-tail, tail-to-head, tail-to-tail) but still incurs quadratic growth, GraphOracle constructs a far sparser Relation-Dependency Graph (RDG) by keeping only directed precedence edges mined from two-hop relational motifs, reducing the edge count to $\mathcal{O}(|\mathcal{R}|\!\cdot\!\bar{d})$. These precedence edges impose an explicit partial order so that information propagates hierarchically from prerequisite to consequent relations, enabling the capture of high-order global dependencies that the local structures of its competitors overlook. Furthermore, while ULTRA directly applies NBFNet’s method for entity and relation representation without specialized relation processing (limiting it to local relation structures), GraphOracle implements a query-conditioned multi-head attention mechanism that traverses the RDG to produce context-specific relation embeddings. This economical yet expressive design suppresses noise, lowers computational cost, and sustains both efficiency and accuracy as the relation set expands, explaining GraphOracle’s consistent superiority in transfer learning across real-world knowledge graphs.
We conducted a controlled experiment to isolate the impact of relation graph construction by replacing GraphOracle’s construction method with those of ULTRA and INGRAM, while maintaining identical message passing mechanisms. The results in Fig. 6 demonstrate that GraphOracle’s relation-dependency graph construction yields consistently superior performance across all datasets and metrics. This empirically validates our theoretical claim that GraphOracle more effectively captures essential compositional relation patterns while filtering out spurious connections that introduce noise into the reasoning process. Notably, while the INGRAM and ULTRA graph construction variants underperform compared to GraphOracle, they still outperform their respective original message passing implementations—further confirming the effectiveness of our attention-based message propagation scheme. The efficiency advantage is quantitatively substantial: as shown in Table 18, GraphOracle generates significantly fewer edges (often 50-60% fewer) than ULTRA and INGRAM across all benchmark datasets. This reduction in graph density translates directly to computational efficiency gains, with GraphOracle requiring proportionally less memory and computation during both training and inference phases. The performance improvements, coupled with this computational efficiency, demonstrate that GraphOracle’s approach to modeling relation dependencies fundamentally addresses the core challenge in knowledge graph foundation models: capturing meaningful compositional patterns without being overwhelmed by the combinatorial explosion of potential relation interactions.